### v1

# Shuai Wang, Xin Zhong

Abstract
We establish the global existence of a class of weak solutions to the isentropic compressible Navier–Stokes equations in a half-plane with Dirichlet boundary conditions, allowing for vacuum both in the interior and at infinity, under a suitably small initial total energy. The solutions constructed here admit unbounded densities and lie in an intermediate regularity regime between the finite-energy weak solutions of Lions–Feireisl and the framework of Hoff. This result generalizes previous works of Hoff (Comm. Pure Appl. Math. 55 (2002), pp. 1365–1407) and Perepelitsa (Arch. Ration. Mech. Anal. 212 (2014), pp. 709–726) concerning discontinuous solutions by allowing vacuum states and unbounded density. Our analysis relies on the Green function method and new estimates involving the specific structure of the equations and the geometry of the half-plane. To the best of our knowledge, this is the first result concerning global weak solutions within Hoff’s framework on an unbounded domain that simultaneously accommodates Dirichlet boundary conditions and far-field vacuum. The intermediate-regularity class developed here may be viewed as a natural extension of Hoff’s theory, precisely tailored to overcome the two corresponding obstructions: the lack of global space-time control of the effective viscous flux arising from far-field vacuum and the absence of boundary-induced regularity gains in the no-slip setting.
Key words and phrases. Compressible Navier–Stokes equations; global weak solutions; vacuum; Green function; Dirichlet boundary conditions.
2020 Mathematics Subject Classification. 35A01; 35B65; 76N10.
1 Introduction
1.1 Background and motivation
We study the isentropic compressible Navier–Stokes equations in the half-plane $\mathbb{R}^{2}_{+}=\{\mathbf{x}\in\mathbb{R}^{2}:x_{2}>0\}$:
$\begin{cases}\rho_{t}+\operatorname{div}(\rho\mathbf{u})=0,\\ (\rho\mathbf{u})_{t}+\operatorname{div}(\rho\mathbf{u}\otimes\mathbf{u})+\nabla P=\mu\Delta\mathbf{u}+(\mu+\lambda)\nabla\operatorname{div}\mathbf{u}\end{cases}$ (1.1)
with the given initial data and far-field conditions
$(\rho,\rho\mathbf{u})|_{t=0}=(\rho_{0},\rho_{0}\mathbf{u}_{0})(\mathbf{x}),~~\mathbf{x}\in\mathbb{R}^{2}_{+},$ (1.2)
$(\rho_{0},\rho_{0}\mathbf{u}_{0})(\mathbf{x})\rightarrow(0,\mathbf{0})\ \ \text{as}\ |\mathbf{x}|\rightarrow\infty,\ \ \mathbf{x}\in\mathbb{R}^{2}_{+},$ (1.3)
and Dirichlet boundary conditions
$\mathbf{u}(\mathbf{x},t)=\mathbf{0},\ \ x_{2}=0,\ t>0,$ (1.4)
where $\mathbf{x}=(x_{1},x_{2})$ is the spatial coordinate, and $t\geq 0$ is the time. The unknowns $\rho\geq 0$, $\mathbf{u}=(u^{1},u^{2})$, and $P=P(\rho)=a\rho^{\gamma}\ (a>0,\gamma>1)$ are the fluid density, velocity, and pressure, respectively. The constants $\mu$ and $\lambda$ represent the shear viscosity and bulk viscosity of the fluid, respectively, satisfying the physical restrictions
$\mu>0,\ \ \ \mu+\lambda\geq 0.$
The global well-posedness of the multi-dimensional isentropic compressible Navier–Stokes equations has been studied extensively, motivated by both the rich mathematical structure of the system and its fundamental physical relevance. In the following, we restrict our discussion to those works that are closely related to the present work. Matsumura and Nishida [26, 27] first obtained global classical solutions for the three-dimensional initial-boundary value problem, under the assumption that the initial data are sufficiently close to constant equilibrium states in $H^{3}$. Later on, Danchin [6] established the global existence of strong solutions for small initial data $(\rho_{0},\mathbf{u}_{0})$ in the scaling-critical Besov spaces $\dot{B}^{\frac{N}{2}}_{2,1}(\mathbb{R}^{N})\times\dot{B}^{\frac{N}{2}-1}_{2,1}(\mathbb{R}^{N})\ (N\geq 2)$. This critical framework was subsequently extended to the $L^{p}$ setting in [4, 5, 11]. It should be noted that all of the above results require the initial density to be strictly positive, thereby excluding the presence of vacuum states.
To advance the theory of compressible flows, increasing attention has been paid to regimes that admit vacuum. In this setting, the degeneracy of the governing equations, together with the evolution of vacuum boundaries, places the problem beyond the scope of standard degenerate PDE theory and necessitates the development of new analytical approaches. The seminal breakthrough is due to P.-L. Lions [25], who established finite-energy weak solutions for the isentropic system with vacuum, satisfying the global energy inequality and characterized by densities in $L^{\infty}_{t}L^{\gamma}_{x}$ with $\gamma\geq\frac{3N}{N+2}$ for $N=2,3$. The key for the construction was to build strong compactness of the density sequence via developing the renormalization techniques and the effective viscous flux arguments. This foundation has inspired significant extensions. Feireisl and collaborators [8, 9] relaxed the restriction on the adiabatic exponent to $\gamma>\frac{N}{2}$ with the aid of oscillation defect measures. Under assumptions of spherical symmetry and axisymmetry, Jiang and Zhang [20, 21] justified global weak solutions for all $\gamma>1$ by deriving refined estimates for the effective viscous flux that lead to improved integrability of the density. Recently, Bresch and Jabin [2] extended the Lions–Feireisl framework to incorporate more complex physics, including non-monotone pressure laws and anisotropic viscous stress tensors. Despite these advances, the question of uniqueness remains a fundamental and largely open problem in the theory of weak solutions.
Parallel to the development of weak solutions theory, substantial progress has been made on the global well-posedness of classical solutions in the presence of vacuum. A significant contribution was made by Huang, Li, and Xin [19], who established the global existence and uniqueness of classical solutions in $\mathbb{R}^{3}$ with vacuum for initial data with small total energy but possibly large oscillations. To address the distinct logarithmic decay at infinity in two dimensions, this framework was extended to the whole plane in [23] through the development of weighted estimates and a refined decay analysis. More recently, a relaxation of the small-energy assumption was achieved in [18] for the Cauchy problem as the adiabatic exponent $\gamma$ approaches $1$. It was shown in [28] that local smooth 3D solutions with $\gamma\leq 1+2/\sqrt{3}$ explode in finite time: the $L^{\infty}$-norms of the density and the velocity blow-up. Thus, in some sense, the smallness condition is necessary in order to ensure global well-posedness. It should be emphasized that one of the key issues in [18, 19, 23] is to derive time-independent uniform upper bounds of the density. For more studies on the compressible Navier–Stokes equations, we refer the reader to the excellent handbook [10] and references contained therein. Let us point out that although the weighted energy method is effective in handling classical solutions as in [23], it is difficult to deal with the theory of global weak solutions in $\mathbb{R}^{2}$ with far-field vacuum due to the lack of decay-sensitive compactness for either the density or the pressure, which obstructs a satisfactory treatment of the nonlinear terms.
Returning to the theory of weak solutions, another significant category is formed by Hoff’s intermediate weak solutions. In his foundational works [12, 13], Hoff justified the global existence of such solutions in the whole space, under the specific structural assumptions that the initial density $\rho_{0}$ is close to a constant in $L^{\infty}$. These solutions possess regularity intermediate between that of Lions–Feireisl’s weak solutions and standard strong solutions. In particular, particle trajectories are well-defined in non-vacuum regions, while density discontinuities may be transported (see [14]). A key feature of Hoff’s approach is the use of a Lagrangian formulation along particle paths, which combined with the $L^{\infty}$ bound on the density simplifies the compactness analysis of both convective and pressure terms. Moreover, although the velocity field is merely log-Lipschitz, the effective viscous flux enjoys Lipschitz regularity. Exploiting this gain, Hoff and Santos [16, 17] established uniqueness and continuous dependence, thereby achieving the global well-posedness theory. Nevertheless, even within this framework, the treatment of far-field vacuum remains a major unresolved challenge.
It should be mentioned that the results discussed above focus primarily on the Cauchy problem, while the initial-boundary value problem for isentropic compressible flows in domains with physical boundary conditions is also mathematically distinct and equally important. For small-energy initial data, Hoff [15] constructed global weak solutions with vacuum in the half-space $\mathbb{R}_{+}^{3}$ under Navier-slip boundary conditions. Global classical solutions with vacuum were later obtained under similar assumptions in the half-space [7] and in bounded smooth domains [3]. A crucial analytical advantage in these settings is that slip boundary conditions provide explicit boundary information for the effective viscous flux $F=(2\mu+\lambda)\operatorname{div}\mathbf{u}-P(\rho)$, a quantity central to the derivation of enhanced regularity. By contrast, despite its physical relevance, progress under no-slip boundary conditions has been much more limited. In this regard, Perepelitsa [31] extended Hoff’s framework to Dirichlet (no-slip) boundaries under the assumptions that the initial density is strictly positive and belongs to a boundary-adapted Hölder space defined by reflection, and that the viscosity ratio $\mu/(\mu+\lambda)$ is suitably small. However, the approach in [31] appears difficult to extend to genuine no-slip boundaries in the absence of additional structural assumptions on the density near the boundary, particularly in the presence of vacuum.
It is worth noting that the propagation of an $L^{\infty}$ bound for the density relies crucially on the algebraic structure of the effective viscous flux (see [14]). To overcome the intrinsic obstacles associated with Dirichlet boundary conditions, it may therefore be necessary to move beyond the classical isentropic framework and to consider regimes in which this structural quantity no longer plays a decisive role. In this spirit, Bresch and Burtea [1] established global weak solutions on the torus $\mathbb{T}^{N}$ for compressible flows with an anisotropic viscous stress tensor, requiring only an $L^{p}$ bound on the density instead of the stronger $L^{\infty}$ condition. Their strategy bypasses the need for regularity gain from the effective viscous flux and thus suggests a potential pathway for addressing the isentropic no-slip problem. On the other hand, the existence of a unique global strong solution in $\mathbb{R}^{3}$ with far-field vacuum was recently obtained in [32] under a smallness assumption on a scaling-invariant quantity, with the analysis relying on the propagation of an additional $L^{2\gamma}$ bound for the density. From a functional analytic perspective, an intermediate $L^{p}$ regularity with $\gamma<p<\infty$ places the pressure $P(\rho)=a\rho^{\gamma}$ in a Lebesgue space that lies strictly between the $L^{1}$ setting of the Lions–Feireisl theory and the $L^{\infty}$ framework of Hoff. This level of spatial integrability provides access to hidden decay information at infinity, which may be instrumental in the treatment of far-field vacuum.
Motivated by these considerations, the present work aims to address these intertwined challenges by constructing global weak solutions within an $L^{p}$ framework for the isentropic compressible Navier–Stokes equations in the half-plane, subject to Dirichlet boundary conditions and allowing vacuum both at interior and infinity. The main difficulty lies not only in each aspect separately, but also in their strong interaction. On the boundary side, several analytical tools, including the effective viscous flux mechanism and vorticity estimates, are no longer available. Meanwhile, the framework of weak solutions provides no effective decay control at infinity, particularly for $\nabla\mathbf{u}$. As a result, one faces a simultaneous loss of boundary structure and far-field decay, which renders many classical approaches inapplicable. To overcome these coupled obstructions, our key innovations are twofold: we mitigate the time growth in the far-field analysis by exploiting the $L^{q}$ $(1<q<\infty$) integrability of the pressure $P(\rho)$, and we decompose the effective viscous flux via the Green function method to obtain the required estimates under the Dirichlet boundary condition.
1.2 Main result
Before stating our main results, we first formulate the notations and conventions used throughout this paper. We denote by $C$ a generic positive constant which may vary at different places, and write $C(f)$ to emphasize its dependence on $f$. The symbol $\Box$ marks the end of a proof and $a\triangleq b$ means $a=b$ by definition. For $1\leq p\leq\infty$ and integer $k\geq 0$, we denote the standard Sobolev spaces:
$L^{p}=L^{p}(\mathbb{R}^{2}_{+}),\ W^{k,p}=W^{k,p}(\mathbb{R}^{2}_{+}),\ H^{k}=W^{k,2},\ D_{0}^{1}=\big\{\mathbf{u}\in L^{1}_{\operatorname{loc}}|\nabla\mathbf{u}\in L^{2},\ \text{the trace}\,\operatorname{tr}(\mathbf{u})=\mathbf{0}\,\ \text{on}\ \partial\mathbb{R}^{2}_{+}\big\}.$
Let $B_{R}$ denote the open ball of radius $R$ centered at the origin. For simplicity, we write
$\widetilde{B}_{R}=B_{R}\cap\mathbb{R}^{2}_{+},\ \ \int f\mathrm{d}\mathbf{x}=\int_{\mathbb{R}^{2}_{+}}f\mathrm{d}\mathbf{x},\ \ f_{i}=\partial_{i}f=\frac{\partial f}{\partial x_{i}}.$
For two $n\times n$ matrices $A=\{a_{ij}\}$ and $B=\{b_{ij}\}$, the symbol $A:B$ represents the trace of the matrix product $AB^{\top}$, i.e.,
$A:B\triangleq\operatorname{tr}(AB^{\top})=\sum_{i,j=1}^{n}a_{ij}b_{ij}.$
The initial total energy is defined as
$C_{0}\triangleq\int_{\mathbb{R}^{2}_{+}}\bigg(\frac{1}{2}\rho_{0}|\mathbf{u}_{0}|^{2}+\frac{1}{\gamma-1}P(\rho_{0})\bigg)\mathrm{d}\mathbf{x}.$ (1.5)
Moreover, we denote by
$\begin{cases}\dot{f}\triangleq f_{t}+\mathbf{u}\cdot\nabla f,\\ F\triangleq(2\mu+\lambda)\operatorname{div}\mathbf{u}-P(\rho),\\ \operatorname{curl}\mathbf{u}\triangleq-\nabla^{\bot}\cdot\mathbf{u}=-\partial_{2}u^{1}+\partial_{1}u^{2},\end{cases}$ (1.6)
which represent the material derivative of $f$, the effective viscous flux, and the vorticity, respectively.
We recall the definition of weak solutions to the problem (1.1)–(1.4) in the sense of [15].
Definition 1.1.
A pair $(\rho,\mathbf{u})$ is said to be a weak solution to the problem (1.1)–(1.4) provided that
$(\rho,\rho\mathbf{u})\in C([0,\infty);H^{-1}(\mathbb{R}^{2}_{+})),\ \ \nabla\mathbf{u}\in L^{2}(\mathbb{R}^{2}_{+}\times(0,\infty))$
with $(\rho,\mathbf{u})|_{t=0}=(\rho_{0},\mathbf{u}_{0})$. Moreover, for any $t_{2}\geq t_{1}\geq 0$ and test function $(\phi,\boldsymbol{\psi})(\mathbf{x},t)\in C^{1}\big(\overline{\mathbb{R}^{2}_{+}}\times[t_{1},t_{2}]\big)$, with uniformly bounded support in $\mathbf{x}$ for $t\in[t_{1},t_{2}]$ and satisfying $\boldsymbol{\psi}|_{\partial\mathbb{R}^{2}_{+}}=0$, the following identities hold [^1] Throughout this paper, we will use the Einstein summation over repeated indices convention.:
$\int_{\mathbb{R}^{2}_{+}}\rho(\mathbf{x},\cdot)\phi(\mathbf{x},\cdot)\mathrm{d}\mathbf{x}\Big|_{t_{1}}^{t_{2}}$ $=\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{2}_{+}}(\rho\phi_{t}+\rho\mathbf{u}\cdot\nabla\phi)\mathrm{d}\mathbf{x}\mathrm{d}t,$
$\int_{\mathbb{R}^{2}_{+}}(\rho\mathbf{u}\cdot\boldsymbol{\psi})(\mathbf{x},\cdot)\mathrm{d}\mathbf{x}\Big|_{t_{1}}^{t_{2}}$ $=\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{2}_{+}}\big(\rho\mathbf{u}\cdot\boldsymbol{\psi}_{t}+\rho u^{i}\mathbf{u}\cdot\partial_{i}\boldsymbol{\psi}+P\operatorname{div}\boldsymbol{\psi}\big)\mathrm{d}\mathbf{x}\mathrm{d}t$
$\quad-\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{2}_{+}}\big(\mu\nabla\mathbf{u}:\nabla\boldsymbol{\psi}+(\mu+\lambda)\operatorname{div}\mathbf{u}\operatorname{div}\boldsymbol{\psi}\big)\mathrm{d}\mathbf{x}\mathrm{d}t.$
Given $\alpha\in(1,2)$, assume that the initial data $(\rho_{0},\mathbf{u}_{0})$ satisfies
$0\leq\rho_{0}\in L^{\theta},\ \ \bar{x}^{\alpha}\rho_{0}\in L^{1},\ \ \mathbf{u}_{0}\in D^{1}_{0},\ \ \rho_{0}|\mathbf{u}_{0}|^{2}\in L^{1},$ (1.7)
where
$\theta\triangleq\frac{12\gamma(2\alpha+1)}{\alpha-1}\in(60\gamma,\infty),\ \ \bar{x}\triangleq\big(e+|\mathbf{x}|^{2}\big)^{1/2}\log^{2}\big(e+|\mathbf{x}|^{2}\big).$ (1.8)
Without loss of generality, let the initial density $\rho_{0}$ verify
$\int_{\mathbb{R}^{2}_{+}}\rho_{0}\mathrm{d}\mathbf{x}=1,$ (1.9)
which implies that there exists a constant $\eta_{0}>0$ such that
$\int_{\widetilde{B}_{\eta_{0}}}\rho_{0}\mathrm{d}\mathbf{x}\geq\frac{1}{2}\int_{\mathbb{R}^{2}_{+}}\rho_{0}\mathrm{d}\mathbf{x}=\frac{1}{2}.$ (1.10)
Moreover, we assume that there exist two constants $\hat{\rho}\geq 1$ and $M\geq 1$ such that
$\|\rho_{0}\|_{L^{\theta}}\leq\hat{\rho},\ \ \|\bar{x}^{\alpha}\rho_{0}\|_{L^{1}}+\|\nabla\mathbf{u}_{0}\|_{L^{2}}\leq M,$ (1.11)
and impose the condition that
$\frac{3\mu}{3\mu+\lambda}\leq\min\bigg\{\frac{1}{2\Lambda(4)},a^{-1}\Lambda^{-1}\bigg(\frac{25\alpha+11}{\alpha-1}\bigg)\bigg\},$ (1.12)
where $\Lambda(\cdot)$ denotes the positive constant introduced in Lemma 2.7.
Now we state our main result on the global existence of weak solutions.
Theorem 1.1.
Let the assumptions (1.7), (1.9), (1.11), and (1.12) be satisfied. Then there exists a positive constant $\varepsilon$ depending only on $\alpha,\hat{\rho},M,a,\gamma,\mu,\lambda$, and $\eta_{0}$ such that if
$C_{0}\leq\varepsilon,$ (1.13)
the problem (1.1)–(1.4) admits a global weak solution $(\rho,\mathbf{u})$ in the sense of Definition 1.1 satisfying, for any $0<T<\infty$,
$\begin{cases}0\leq\rho\in L^{\infty}(0,T;L^{\theta}(\mathbb{R}^{2}_{+}))\cap C([0,T];L^{q}(\mathbb{R}^{2}_{+})),\ \ \text{for any}\ q\in[1,\theta),\\ \bar{x}^{\alpha}\rho\in L^{\infty}(0,T;L^{1}(\mathbb{R}^{2}_{+})),\ \ \rho\mathbf{u}\in C([0,T];H^{-1}(\mathbb{R}^{2}_{+})),\\ \nabla\mathbf{u}\in L^{2}(\mathbb{R}^{2}_{+}\times(0,T)),\ \ \big(\sigma^{\frac{1}{2}}\sqrt{\rho}\dot{\mathbf{u}},\ \sigma^{\frac{3}{2}}\nabla\dot{\mathbf{u}}\big)\in L^{2}(\mathbb{R}^{2}_{+}\times(0,T)),\end{cases}$ (1.14)
with $\sigma=\sigma(t)\triangleq\min\{1,t\}$, and
$\inf_{0\leq t\leq T}\int_{\widetilde{B}_{\eta_{1}(1+t)\log^{\alpha}(e+t)}}\rho(\mathbf{x},t)\mathrm{d}\mathbf{x}\geq\frac{1}{4}$ (1.15)
for some positive constant $\eta_{1}$ depending only on $\alpha,\hat{\rho},M,a,\gamma,\eta_{0}$, and $\|\sqrt{\rho_{0}}\mathbf{u}_{0}\|_{L^{2}}$.
Several remarks are in order.
Remark 1.1.
It should be noted that the weak solutions constructed in Theorem 1.1 occupy an intermediate regularity regime between Lions–Feireisl’s finite-energy weak solutions and the weak solutions developed by Hoff [12, 13, 14, 15]. Our framework of a half-plane with far-field vacuum and Dirichlet boundary conditions provides significant insight for two extremes: the Cauchy problem with vacuum at infinity and the Dirichlet problem in bounded domains. We leave such interesting problems for our future studies.
Remark 1.2.
The present work improves the result of Perepelitsa [31]. On the one hand, the presence of vacuum is allowed both at interior and infinity. On the other hand, in contrast to the $L^{\infty}$ density assumption in [31], our solutions are constructed within an $L^{\theta}$ density framework, which permits unbounded densities and enables a reduction of the time growth in the far-field analysis through the relaxed spatial integrability of the pressure in the sense that $P(\rho)\in L^{q}$ for $1<q=\theta/\gamma<\infty$.
Remark 1.3.
It should be emphasized that the approach introduced by Bresch and Burtea [1] relies essentially on the periodic geometry of the torus $\mathbb{T}^{N}$, which allows them to make use of harmonic analysis tools due to the absence of boundaries, and therefore cannot be directly extended to the setting with physical boundary conditions. The absence of an $L^{\infty}$ bound for the density constitutes a fundamental analytical difference in the treatment of the nonlinear terms; see Subsection 1.3 for details. Conversely, the intermediate-regularity viewpoint adopted here may serve as a useful guideline for more complex models, including compressible non-Newtonian flows, where pointwise density bounds are out of reach and integrability-based structural estimates become essential.
Remark 1.4.
The index $\theta$ in (1.8) is a technical parameter determined by weighted inequalities in Lemma 2.4 and large-time analysis in (3), and may be relaxed by delicate analysis; see Subsection 1.3 for details. In addition, it should be mentioned that it is not clear whether the regularity assumption on the initial velocity in Theorem 1.1 can be generalized to $H^{s}$ for some $0<s<1$, which has been achieved in the whole plane case [14] as the initial density is bounded from above and below (among other conditions).
Remark 1.5.
The present framework is well suited to the treatment of far-field vacuum, but it does not readily accommodate non-vacuum far-field states. Indeed, the key estimate
$\|\rho\dot{\mathbf{u}}\|_{L^{p}}\leq\|\sqrt{\rho}\|_{L^{\frac{2p}{2-p}}}\,\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}},\qquad 1<p<2,$
cannot be effectively exploited without sufficient decay of the density at infinity. In two dimensions, this inequality, combined with the $L^{q}$-integrability of the pressure (taken as $q=60$ in this work), yields the improved time-growth bounds used in the far-field analysis; see Subsection 1.3 for details.
As a key ingredient in the treatment of the Dirichlet boundary condition, we employ the Green function method together with the boundary-compatible quantity $\mathbf{x}F$, which satisfies
$x_{2}F=0\quad\text{on }\partial\mathbb{R}^{2}_{+}=\{x_{2}=0\}.$
This observation is specific to the half-plane geometry; see Subsection 2.2 for details. Motivated by this structure, it would be natural to investigate whether an analogous boundary-compatible mechanism can be identified in bounded domains, for instance in the unit ball, by considering a suitable condition involving $(|\mathbf{x}|^{2}-1)F$ on the boundary.
Finally, we note that in three dimensions substantially new ideas seem necessary, since new phenomena arise and several crucial estimates break down, including the above inequality and the control of $\|\nabla\mathbf{u}\|_{L^{6}}$; see (2.2). Addressing such problems requires new observations and techniques, which are left to investigate in future.
1.3 Strategy of the proof
We now outline the main ideas and difficulties in the proof. The argument relies on the construction of global smooth approximate solutions via a delicate limiting procedure. This approach consists of two fundamental steps. First, we apply the local existence theory for initial densities that are strictly positive (Lemma 2.1). Second, we pass to the limit as the lower bound of the initial density tends to zero (see Section 4 for details). Bridging these two steps crucially depends on deriving suitable a priori estimates that are uniform with respect to the lower bound of the initial density and ensure the desired regularity.
It should be emphasized that the key techniques developed in [15, 1] do not appear to be directly adaptable to the Dirichlet boundary value problem. In addition, without an $L^{\infty}$ bound of the density, it is unclear how the approach of [31] could be extended, particularly in the presence of vacuum both in the interior and at spatial infinity. As a consequence, new observations and ideas are required to handle the joint presence of no-slip boundary conditions and far-field vacuum.
[…]
$\frac{\mathrm{d}}{\mathrm{d}t}\int\rho^{r}\mathrm{d}\mathbf{x}+(r-1)\int\rho^{r}\operatorname{div}\mathbf{u}\mathrm{d}\mathbf{x}=0$ (1.16)
for any $1<r<\infty$. This indicates that the effective viscous flux
$F=(2\mu+\lambda)\operatorname{div}\mathbf{u}-P(\rho)$
plays a crucial role in generating dissipation for the density at the $L^{r}$ level. However, the presence of the Dirichlet boundary necessitates a substantially more delicate analysis of both $F$ and $P(\rho)$, especially under the assumption that $\rho\in L^{1}\cap L^{\theta}$.
The first and foremost difficulty lies in the treatment of the effective viscous flux $F$, which therefore demands and indeed receives substantial analytical attention throughout the proof (see Subsection 2.2 for details). Unlike the cases of $\mathbb{R}^{N}$ or bounded domains with slip boundary conditions, the Dirichlet setting yields weaker analytic control from the structural quantity for enhancing regularity. This in turn motivates the introduction of a quantity that formally shares the essential structural advantages of Hoff’s framework. Relying solely on the equations and boundary conditions, we observe that the specific geometry of the half-plane gives rise to an additional boundary
$x_{2}F=-F\mathbf{x}\cdot\mathbf{n}=0,\ \ \text{on}\ \partial\mathbb{R}^{2}_{+}=\{x_{2}=0\},$
where $\mathbf{n}=(0,-1)$ is the unit outward normal vector to $\partial\mathbb{R}^{2}_{+}$. Based on this observation, we exploit the elliptic structures of $\Delta\mathbf{u}$ and $\Delta(\mathbf{x}F)$. In combination with the Green function method, this allows us to decompose $F$ into two components (see Lemma 2.6 for details):
$F\sim\widetilde{F}+\mathcal{Q}_{3}(P),$
where (see Lemma 2.7)
$\|\nabla\widetilde{F}\|_{L^{q}}\lesssim\|\rho\dot{\mathbf{u}}\|_{L^{q}},\ \ \|\mathcal{Q}_{3}(P)\|_{L^{q}}\lesssim\|P\|_{L^{q}},\ \ \ \ \ 1<q<\infty.$
Once the $L^{q}$ bound for the pressure $P$ is established, the analysis effectively reduces to studying $\widetilde{F}$, which serves as a more regular and analytically tractable principal unknown.
The second major difficulty arises in obtaining estimates for $\|\nabla\mathbf{u}\|_{L^{q}}$ with $1<q<\infty$. In this setting, the classical approach based on the Helmholtz decomposition is no longer applicable, due to the absence of an $L^{q}$ bound for the vorticity $\operatorname{curl}\mathbf{u}$. Indeed, although one may isolate a certain component from $F$, the remaining part, when coupled with $\operatorname{curl}\mathbf{u}$, fails to provide any additional regularity. As an alternative strategy, we decompose the velocity field by analyzing separately the contributions of the source terms in the Lamé system (see (2.20) for $\mathbf{u}=\mathbf{w}_{1}+\mathbf{w}_{2}$). This approach yields a cruder estimate for $\nabla\mathbf{u}$, which forces us to go beyond the framework of $L^{4}$ estimates and to employ $L^{6}$ bounds at certain stages of the analysis (see (3) and (3)). In light of this, we adopt the following estimate to the two-dimensional setting (see (2.2)):
$\|\nabla\mathbf{w}_{1}\|_{L^{6}}\lesssim\|\nabla^{2}\mathbf{w}_{1}\|_{L^{3/2}}\lesssim\|\rho\dot{\mathbf{u}}\|_{L^{3/2}}\lesssim\|\sqrt{\rho}\|_{L^{6}}\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}.$
Moreover, obtaining any decay control at infinity within the weak solution framework is particularly delicate and challenging, especially in the presence of Dirichlet boundary conditions. This is due to the weakness of the available a priori bounds, such as the crude control of $\|\nabla\mathbf{u}\|_{L^{q}}$, and to the lack of higher-order estimates, which leaves no mechanism to derive decay in this setting. In view of the tools at hand, the only information we can employ is the $L^{\theta}$ bound concerning the density. As an attempt, we exploit the material derivative weighted by the density (see (3.56) and (3.59)), and obtain only a very poor growth bound:
$\|\nabla\widetilde{F}\|_{L^{4}}\lesssim\|\rho\dot{\mathbf{u}}\|_{L^{4}}\lesssim(1+t)^{4}.$
Unlike in the strong-solution setting, the route based on the decay-control for higher derivatives of the velocity field is no longer available, so the decay mechanism in our setting leads us to re-examine the density and the pressure $P(\rho)$. By exploiting the higher $L^{60}$ integrability of the pressure together with the spatial embedding in two-dimensional setting, we sharpen the above estimate and reduce the time growth (see (3)):
$\|\widetilde{F}\|_{L^{\infty}}\lesssim\|\widetilde{F}\|_{L^{60}}^{\frac{15}{16}}\|\nabla\widetilde{F}\|_{L^{4}}^{\frac{1}{16}}\lesssim(1+t)^{\frac{1}{4}},$
which can now be enclosed. Furthermore, this integrability requirement together with Lemma 2.4 dictates our choice of the index
$\theta=\max\big\{4(2\alpha+1)/(\alpha-1),60\gamma\big\}.$
As a heuristic indication for a possible refinement of this index, we reconsider the growth bounds (2.1) and (3.56) with $\|\rho\dot{\mathbf{u}}\|_{L^{r}}\lesssim(1+t)^{4}$ for $2<r<\infty$. To accommodate the far-field behavior, the choice of the index is guided by the Gagliardo–Nirenberg inequality (see Lemma 2.2) with
$\|\widetilde{F}\|_{L^{q}}^{\frac{q(r-2)}{2r+q(r-2)}}\lesssim\Big(\|\rho\dot{\mathbf{u}}\|_{L^{\frac{2q}{2+q}}}+\|P\|_{L^{q}}\Big)^{\frac{q(r-2)}{2r+q(r-2)}},\ \ \ \|\nabla\widetilde{F}\|_{L^{r}}^{\frac{2r}{2r+q(r-2)}}\lesssim\|\rho\dot{\mathbf{u}}\|_{L^{r}}^{\frac{2r}{2r+q(r-2)}}\lesssim(1+t)^{\frac{1}{4}},\ \ q=\frac{30r}{r-2},$
owing to the decomposition of the velocity field in Lemma 2.8. Accordingly, the refined initial density integrability index $\tilde{\theta}$ and the auxiliary exponent $\tilde{\zeta}(\alpha,r)$ in Lemma 2.3 are dictated by
$\tilde{\theta}=\max\big\{\tilde{\zeta}(\alpha,r),q\gamma=30\gamma r/(r-2)\big\}>30\gamma,$
where we take $r=4$ in the present work.
Finally, the $L^{\theta}$ integrability of the density leads to additional analytical difficulties in the treatment of the nonlinear terms involving $P(\rho)$, thereby necessitating a more refined analysis. For example, the term
$\int_{0}^{T}\int\sigma P|\nabla\mathbf{u}|^{2}\mathrm{d}\mathbf{x}\mathrm{d}t$
becomes difficult to control analytically for $\sigma=\sigma(t)\triangleq\min\{1,t\}$, particularly in view of the lack of a satisfactory bound for $\|\nabla\mathbf{u}\|_{L^{4}}$. To overcome this obstacle, we split the time interval into two regimes and carefully isolate the singular behavior of the pressure term near the initial time (see (3) and (3.20)). This procedure allows us to focus on
$\int_{0}^{T}\int\sigma^{3}P^{4}\mathrm{d}\mathbf{x}\mathrm{d}t,$
which can be enclosed by the estimates derived in (3). In a certain sense, the interplay between reduced density regularity and the Dirichlet boundary condition manifests itself in stronger temporal singularities near the initial time.
Having resolved the above difficulties, we further apply Zlotnik’s inequality (see Lemma 2.5) to obtain a time-uniform upper bound for $\|\rho\|_{L^{\theta}}$. A key step in this procedure is the control of $\|\widetilde{F}\|_{L^{\infty}}$, as shown in (3) and (3), where the density-weighted inequality established in Lemma 2.4 and the spatial weighted estimate in Lemma 3.5 play central roles. With these a priori estimates in hand, we obtain the desired regularity for the density, which is higher than that of Lions–Feireisl’s class. At last, by a standard compactness argument combined with the continuity method, we construct a global weak solution in the sense of Definition 1.1 that satisfies (1.14).
The rest of the paper is organized as follows. In the next section, we recall some known facts and elementary inequalities to be used later, and we introduce the core treatment of the effective viscous flux $F$ in Subsection 2.2. Section 3 is devoted to establishing a priori estimates. The proof of Theorem 1.1 is given in Section 4.
2 Preliminaries
In this section we collect some facts and elementary inequalities that will be used later.
2.1 Auxiliary results and inequalities
In this subsection we review some known facts and inequalities. First, similar to the proofs in [26, 22], we have the following result concerning the local existence of strong solutions to the problem (1.1)–(1.4).
Lemma 2.1.
Assume that
$\bar{x}^{\alpha}\rho_{0}\in L^{1},\ \rho_{0}\in H^{2},\ \mathbf{u}_{0}\in D^{2}\cap D_{0}^{1},\ \sqrt{\rho_{0}}\mathbf{u}_{0}\in L^{2},\ \inf\limits_{\mathbf{x}\in\widetilde{B}_{R}}\rho_{0}(\mathbf{x})>0$
for some large half-ball $\widetilde{B}_{R}$, then there exists a small time $T$ such that the problem (1.1)–(1.4) admits a unique strong solution $(\rho,\mathbf{u})$ satisfying
$\bar{x}^{\alpha}\rho\in L^{\infty}(0,T;L^{1}),\ \ \rho\in C([0,T];H^{2}),\ \mathbf{u}\in C([0,T];D^{2}\cap D^{1}),\ \inf_{\widetilde{B}_{R}\times[0,T]}\rho(\mathbf{x},t)\geq\frac{1}{2}\inf\limits_{\mathbf{x}\in\widetilde{B}_{R}}\rho_{0}(\mathbf{x})>0.$
The following well-known Gagliardo–Nirenberg inequality (see [29]) will be used later.
Lemma 2.2.
For $p\in[2,\infty)$, $q\in(1,\infty)$, and $r\in(2,\infty)$, there exists some generic constant $C>0$ which may depend on $p$, $q$, and $r$ such that, for $f\in H^{1}$ and $g\in L^{q}$ with $\nabla g\in L^{r}$,
$\|f\|_{L^{p}}\leq C\|f\|_{L^{2}}^{\frac{2}{p}}\|\nabla f\|_{L^{2}}^{1-\frac{2}{p}},\ \ \ \|g\|_{L^{\infty}}\leq C\|g\|_{L^{q}}^{\frac{q(r-2)}{2r+q(r-2)}}\|\nabla g\|_{L^{r}}^{\frac{2r}{2r+q(r-2)}}.$
For $\mathbf{v}\in H^{1}_{0}(\mathbb{R}^{2}_{+})$, define its odd extension $\tilde{\mathbf{v}}$ to $\mathbb{R}^{2}$. Then $\tilde{\mathbf{v}}\in H^{1}(\mathbb{R}^{2})$ and there exists an absolute constant $C>0$ such that
$C^{-1}\|\mathbf{v}\|_{H^{1}(\mathbb{R}^{2}_{+})}\leq\|\tilde{\mathbf{v}}\|_{H^{1}(\mathbb{R}^{2})}\leq C\|\mathbf{v}\|_{H^{1}(\mathbb{R}^{2}_{+})}.$
Thus, via this extension argument, the following weighted inequality in [24, Theorem B.1], originally stated for $\mathbb{R}^{2}$, also holds for functions in $H^{1}_{0}(\mathbb{R}^{2}_{+})$.
Lemma 2.3.
For $m\in[2,\infty)$ and $\beta\in(1+m/2,\infty)$, there exists a positive constant $C$ such that for all $\mathbf{v}\in H^{1}_{0}(\mathbb{R}^{2}_{+})$,
$\left(\int_{\mathbb{R}^{2}_{+}}\frac{|\mathbf{v}|^{m}}{e+|\mathbf{x}|^{2}}(\log(e+|\mathbf{x}|^{2}))^{-\beta}\mathrm{d}\mathbf{x}\right)^{1/m}\leq C\|\mathbf{v}\|_{L^{2}(\widetilde{B}_{1})}+C\|\nabla\mathbf{v}\|_{L^{2}(\mathbb{R}^{2}_{+})}.$
Based on Lemma 2.3 and Poincaré’s inequality, we obtain the key bound for $\|\varrho\mathbf{v}\|_{L^{4}}$ with a nonnegative function $\varrho\in L^{1}\cap L^{\zeta}$ ($\zeta<\theta$), which extends the inequality in [23, Lemma 2.4] with $\varrho\in L^{1}(\mathbb{R}^{2})\cap L^{\infty}(\mathbb{R}^{2})$.
Lemma 2.4.
Let $\zeta=4(2\alpha+1)/(\alpha-1)$ with $\alpha\in(1,\infty)$ and $\bar{x}$ be as in (1.8). Assume that the nonnegative function $\varrho\in L^{1}(\mathbb{R}^{2}_{+})\cap L^{\zeta}(\mathbb{R}^{2}_{+})$ satisfies
$M_{2}\leq\int_{\widetilde{B}_{\eta_{*}}}\varrho\mathrm{d}\mathbf{x}\leq\|\varrho\|_{L^{1}(\mathbb{R}^{2}_{+})}=M_{1},\ \ \|\varrho\|_{L^{\zeta}(\mathbb{R}^{2}_{+})}\leq M_{3},\ \ \bar{x}^{\alpha}\varrho\in L^{1}(\mathbb{R}^{2}_{+})$
for positive constants $M_{1},M_{2},M_{3}$, and $\eta_{*}\geq 1$, where $\widetilde{B}_{\eta_{*}}=B_{\eta_{*}}\cap\mathbb{R}^{2}_{+}$. Then there exists a constant $C>0$ depending only on $M_{1}$, $M_{2}$, and $M_{3}$ such that, for $\mathbf{v}\in D^{1}_{0}(\mathbb{R}^{2}_{+})$ verifying $\sqrt{\varrho}\mathbf{v}\in L^{2}(\mathbb{R}^{2}_{+})$,
$\|\varrho\mathbf{v}\|_{L^{4}(\mathbb{R}^{2}_{+})}^{4}\leq C\eta_{*}^{8}(1+\|\bar{x}^{\alpha}\varrho\|_{L^{1}(\mathbb{R}^{2}_{+})})\big(\|\sqrt{\varrho}\mathbf{v}\|_{L^{2}(\mathbb{R}^{2}_{+})}^{4}+\|\nabla\mathbf{v}\|_{L^{2}(\mathbb{R}^{2}_{+})}^{4}\big).$ (2.1)
Proof.
Define the average of $f$ over the domain $\Omega\subset\mathbb{R}^{2}_{+}$ as
$\langle f\rangle_{\Omega}\triangleq\frac{1}{|\Omega|}\int_{\Omega}f(\mathbf{x})\mathrm{d}\mathbf{x}.$
Using the Poincaré inequality in the zero-average form, we deduce from the odd extension argument that
$\|\mathbf{v}-\langle\mathbf{v}\rangle_{\widetilde{B}_{\eta_{*}}}\|_{L^{2}(\widetilde{B}_{\eta_{*}})}\leq C\eta_{*}\|\nabla\mathbf{v}\|_{L^{2}(\widetilde{B}_{\eta_{*}})},$
which along with Hölder’s inequality yields that
$\big|\langle\varrho\rangle_{\widetilde{B}_{\eta_{*}}}\langle\mathbf{v}\rangle_{\widetilde{B}_{\eta_{*}}}\big|$ $=\left|\frac{1}{|\widetilde{B}_{\eta_{*}}|}\int_{\widetilde{B}_{\eta_{*}}}\big(\langle\varrho\rangle_{\widetilde{B}_{\eta_{*}}}-\varrho\big)\big(\mathbf{v}-\langle\mathbf{v}\rangle_{\widetilde{B}_{\eta_{*}}}\big)\mathrm{d}\mathbf{x}+\frac{1}{|\widetilde{B}_{\eta_{*}}|}\int_{\widetilde{B}_{\eta_{*}}}\varrho\mathbf{v}\mathrm{d}\mathbf{x}\right|$
$\leq C\eta_{*}^{-2}\big(\|\varrho\|_{L^{2}(\widetilde{B}_{\eta_{*}})}+\eta_{*}^{-1}M_{1}\big)\|\mathbf{v}-\langle\mathbf{v}\rangle_{\widetilde{B}_{\eta_{*}}}\|_{L^{2}(\widetilde{B}_{\eta_{*}})}+C\eta_{*}^{-2}\|\sqrt{\varrho}\|_{L^{2}(\widetilde{B}_{\eta_{*}})}\|\sqrt{\varrho}\mathbf{v}\|_{L^{2}(\widetilde{B}_{\eta_{*}})}$
$\leq C\eta_{*}^{-1}\big(\|\varrho\|_{L^{2}(\widetilde{B}_{\eta_{*}})}+M_{1}\big)\|\nabla\mathbf{v}\|_{L^{2}(\widetilde{B}_{\eta_{*}})}+C\eta_{*}^{-2}M_{1}^{1/2}\|\sqrt{\varrho}\mathbf{v}\|_{L^{2}(\widetilde{B}_{\eta_{*}})}.$
Hence, we have
$\big|\langle\mathbf{v}\rangle_{\widetilde{B}_{\eta_{*}}}\big|\leq C(M_{1},M_{2})\eta_{*}\big(1+\|\varrho\|_{L^{2}(\widetilde{B}_{\eta_{*}})}\big)\|\nabla\mathbf{v}\|_{L^{2}(\widetilde{B}_{\eta_{*}})}+C(M_{1},M_{2})\|\sqrt{\varrho}\mathbf{v}\|_{L^{2}(\widetilde{B}_{\eta_{*}})},$
which implies that
$\|\mathbf{v}\|_{L^{2}(\widetilde{B}_{\eta_{*}})}^{2}$ $\leq 2\int_{\widetilde{B}_{\eta_{*}}}\big|\mathbf{v}-\langle\mathbf{v}\rangle_{\widetilde{B}_{\eta_{*}}}\big|^{2}\mathrm{d}\mathbf{x}+2\big|\widetilde{B}_{\eta_{*}}\big|\big|\langle\mathbf{v}\rangle_{\widetilde{B}_{\eta_{*}}}\big|^{2}$
$\leq C(M_{1},M_{2})\eta_{*}^{4}\big(1+\|\varrho\|_{L^{2}(\widetilde{B}_{\eta_{*}})}^{2}\big)\|\nabla\mathbf{v}\|_{L^{2}(\widetilde{B}_{\eta_{*}})}^{2}+C(M_{1},M_{2})\eta_{*}^{2}\|\sqrt{\varrho}\mathbf{v}\|_{L^{2}(\widetilde{B}_{\eta_{*}})}^{2}.$ (2.2)
The local estimate on a half-ball together with Lemma 2.3 leads to
$\|\varrho\mathbf{v}\|_{L^{4}(\mathbb{R}^{2}_{+})}^{4}$ $\leq\|(\bar{x}^{\alpha}\varrho)^{\frac{2}{\alpha+1}}\|_{L^{\frac{\alpha+1}{2}}(\mathbb{R}^{2}_{+})}\|\bar{x}^{-\frac{2\alpha}{\alpha+1}}|\mathbf{v}|^{4}\|_{L^{\frac{2(\alpha+1)}{\alpha-1}}(\mathbb{R}^{2}_{+})}\|\varrho^{\frac{\alpha-1}{\alpha+1}}\|_{L^{\frac{4(\alpha+1)(2\alpha+1)}{(\alpha-1)^{2}}}(\mathbb{R}^{2}_{+})}\|\varrho^{3}\|_{L^{\frac{4(2\alpha+1)}{3(\alpha-1)}}(\mathbb{R}^{2}_{+})}$
$\leq C\eta_{*}^{8}(1+\|\bar{x}^{\alpha}\varrho\|_{L^{1}(\mathbb{R}^{2}_{+})})(1+\|\varrho\|_{L^{2}(\mathbb{R}^{2}_{+})}^{4})\|\varrho\|_{L^{\frac{4(2\alpha+1)}{\alpha-1}}(\mathbb{R}^{2}_{+})}^{3+\frac{\alpha-1}{\alpha+1}}\big(\|\sqrt{\varrho}\mathbf{v}\|_{L^{2}(\mathbb{R}^{2}_{+})}+\|\nabla\mathbf{v}\|_{L^{2}(\mathbb{R}^{2}_{+})}\big)^{4}$
$\leq C(M_{1},M_{2},M_{3})\eta_{*}^{8}(1+\|\bar{x}^{\alpha}\varrho\|_{L^{1}(\mathbb{R}^{2}_{+})})\big(\|\sqrt{\varrho}\mathbf{v}\|_{L^{2}(\mathbb{R}^{2}_{+})}^{4}+\|\nabla\mathbf{v}\|_{L^{2}(\mathbb{R}^{2}_{+})}^{4}\big),$ (2.3)
where we have chosen
$\beta=2\times\frac{4\alpha}{\alpha-1}>1+\frac{4(\alpha+1)}{\alpha-1}.$
Finally, we introduce a variant of Zlotnik’s inequality (see [33]) to derive the uniform (in time) upper bound of $\|\rho\|_{L^{\theta}}$.
Lemma 2.5.
Let the function $y$ satisfy [^2] Although [33, Lemma 1.3(b)] states the result for $y^{\prime}(t)=g(y)+b^{\prime}(t)$, its proof actually works for the inequality as well.
$y^{\prime}(t)\leq g(y)+b^{\prime}(t)\ \ \text{on}\ [0,T],\quad y(0)=y^{0},$
with $g\in C(\mathbb{R})$ and $y,b\in W^{1,1}(0,T)$. If $g(\infty)=$-$\infty$ and
$b(t_{2})-b(t_{1})\leq N_{0}+N_{1}(t_{2}-t_{1})$
for all $0\leq t_{1}<t_{2}\leq T$ with some $N_{0}\geq 0$ and $N_{1}\geq 0$, then
$y(t)\leq\max\{y^{0},\xi^{*}\}+N_{0}<\infty\ \ \text{on}\ [0,T],$
where $\xi^{*}$ is a constant such that
$g(\xi)\leq-N_{1}\ \ \text{for}\ \xi\geq\xi^{*}.$
2.2 An exposition on the boundary
In this subsection we explore the boundary condition in more detail and provide partial $L^{p}$-estimates for the effective viscous flux $F$ and $\nabla\mathbf{u}$.
[…]
$\rho\dot{\mathbf{u}}=\nabla F-\mu\nabla^{\bot}\operatorname{curl}\mathbf{u}.$ (2.4)
In particular, if we let
$(u^{1},u^{2})=(u^{1},0),\ \ \text{on}\ \partial\mathbb{R}^{2}_{+},$
then $\dot{u}^{2}=0$ on $\partial\mathbb{R}^{2}_{+}$, which along with (2.4) implies that
$0=\rho\dot{\mathbf{u}}\cdot\mathbf{n}=\mathbf{n}\cdot(\nabla F-\mu\nabla^{\bot}\operatorname{curl}\mathbf{u})=\mathbf{n}\cdot\nabla F+\mu\partial_{1}\partial_{2}u^{1},\ \ \text{on}\ \partial\mathbb{R}^{2}_{+},$ (2.5)
where $\mathbf{n}=(0,-1)$ is the unit outward normal vector to $\partial\mathbb{R}^{2}_{+}$. Indeed, the Navier-slip boundary conditions $(u^{1},u^{2})|_{x_{2}=0}=(\partial_{2}u^{1},0)|_{x_{2}=0}$ (see [15]) supply the boundary information needed to exploit the elliptic structure of $F$.
By contrast, the Dirichlet boundary conditions (1.4) do not seem to be of use here, and we cannot directly apply the standard elliptic theory for $F$. Fortunately, although no suitable boundary condition exists, the particular feature of the flat boundary suggests that
$x_{2}F=-F\mathbf{x}\cdot\mathbf{n}=0,\ \ \text{on}\ \partial\mathbb{R}^{2}_{+}=\{x_{2}=0\}.$
This observation, along with the boundary conditions (1.4), directs our attention to $\Delta\mathbf{u}$ and $\Delta(\mathbf{x}F)$ instead. Moreover, we need to treat boundary effects separately, for which the Green function method presents itself as a natural choice.
[…]
$\begin{cases}\mu\Delta\mathbf{u}=\rho\dot{\mathbf{u}}-\frac{\mu+\lambda}{2\mu+\lambda}\nabla F+\frac{\mu}{2\mu+\lambda}\nabla P,&\mathbf{x}\in\mathbb{R}^{2}_{+},\\ \mathbf{u}=0,&\mathbf{x}\in\partial\mathbb{R}^{2}_{+}.\end{cases}$ (2.6)
[…]
$\mathbf{v}(\mathbf{x})=\int\mathbb{G}(\mathbf{x},\mathbf{y})\cdot\Delta\mathbf{v}(\mathbf{y})\mathrm{d}\mathbf{y}+\int_{\partial\mathbb{R}^{2}_{+}}\frac{\partial\mathbb{G}(\mathbf{x},\mathbf{y})}{\partial\mathbf{n}_{\mathbf{y}}}\cdot\mathbf{v}(\mathbf{y})\mathrm{d}s\triangleq\mathbb{G}_{0}(\Delta\mathbf{v})+\mathbb{G}_{b}(\mathbf{v}).$ (2.7)
Correspondingly, for a scalar function $h$, we have
$h(\mathbf{x})=\int\widetilde{G}(\mathbf{x},\mathbf{y})\Delta h(\mathbf{y})\mathrm{d}\mathbf{y}+\int_{\partial\mathbb{R}^{2}_{+}}\frac{\partial\widetilde{G}(\mathbf{x},\mathbf{y})}{\partial\mathbf{n}_{\mathbf{y}}}h(\mathbf{y})\mathrm{d}s\triangleq G_{0}(\Delta h)+G_{b}(h).$ (2.8)
Next, using the Green function representation, we present a decomposition of $F$.
Lemma 2.6.
Let $(\rho,\mathbf{u})$ be a smooth solution to the problem (1.1)–(1.4). For the effective viscous flux $F$ as in (1.6), it holds that
$F=\widetilde{F}+\frac{2\mu}{3\mu+\lambda}\mathcal{Q}_{3}\triangleq\Big[\mathcal{Q}_{1}(\rho\dot{\mathbf{u}})+\frac{\mu+\lambda}{3\mu+\lambda}\mathcal{Q}_{2}(\rho\dot{\mathbf{u}})\Big]+\frac{2\mu}{3\mu+\lambda}\mathcal{Q}_{3}(P),$ (2.9)
where
$\begin{cases}\mathcal{Q}_{1}=\mathcal{Q}_{1}(\rho\dot{\mathbf{u}})\triangleq\operatorname{div}\mathbb{G}_{0}(\rho\dot{\mathbf{u}}),\\ \mathcal{Q}_{2}=\mathcal{Q}_{2}(\rho\dot{\mathbf{u}})\triangleq\operatorname{div}\mathbb{G}_{0}(\mathbf{x}\operatorname{div}(\rho\dot{\mathbf{u}}))-\mathbf{x}\cdot\nabla G_{0}(\operatorname{div}(\rho\dot{\mathbf{u}}))+\operatorname{div}\mathbb{G}_{0}(\rho\dot{\mathbf{u}})-G_{0}(\operatorname{div}(\rho\dot{\mathbf{u}})),\\ \mathcal{Q}_{3}=\,\mathcal{Q}_{3}(P)\,\triangleq\operatorname{div}\mathbb{G}_{0}(\nabla P)-P.\end{cases}$
Proof.
Applying operator $\operatorname{div}$ to the system (2.4) yields
$\Delta F=\operatorname{div}(\rho\dot{\mathbf{u}}),$ (2.10)
which combined with (2.8) shows that
$F=G_{0}(\operatorname{div}(\rho\dot{\mathbf{u}}))+G_{b}(F).$ (2.11)
Note that $\Delta(\mathbf{x}F)=\mathbf{x}\Delta F+2\nabla F$. Putting this expression into (2.6) leads to
$\frac{\mu+\lambda}{2(2\mu+\lambda)}\Delta(\mathbf{x}F)+\mu\Delta\mathbf{u}$ $=\frac{\mu+\lambda}{2(2\mu+\lambda)}\mathbf{x}\Delta F+\frac{\mu+\lambda}{2\mu+\lambda}\nabla F+\mu\Delta\mathbf{u}$
$=\frac{\mu+\lambda}{2(2\mu+\lambda)}\mathbf{x}\operatorname{div}(\rho\dot{\mathbf{u}})+\frac{\mu}{2\mu+\lambda}\nabla P+\rho\dot{\mathbf{u}}\triangleq U.$ (2.12)
Using (2.7), one deduces that
$\frac{\mu+\lambda}{2(2\mu+\lambda)}\mathbf{x}F+\mu\mathbf{u}=\mathbb{G}_{0}(U)+\frac{\mu+\lambda}{2(2\mu+\lambda)}\mathbb{G}_{b}(\mathbf{x}F).$
Taking operator $\operatorname{div}$ on both sides of the above equality, we obtain from (2.11) that
$\operatorname{div}\mathbb{G}_{0}(U)+\frac{\mu+\lambda}{2(2\mu+\lambda)}\operatorname{div}\mathbb{G}_{b}(\mathbf{x}F)$
$=\frac{\mu+\lambda}{2(2\mu+\lambda)}(2F+\mathbf{x}\cdot\nabla F)+\mu\operatorname{div}\mathbf{u}$
$=\frac{\mu+\lambda}{2\mu+\lambda}F+\frac{\mu+\lambda}{2(2\mu+\lambda)}\mathbf{x}\cdot\nabla[G_{0}(\operatorname{div}(\rho\dot{\mathbf{u}}))+G_{b}(F)]+\frac{\mu}{2\mu+\lambda}(F+P).$
Thus,
$F+\frac{\mu}{2\mu+\lambda}P=\operatorname{div}\mathbb{G}_{0}(U)-\frac{\mu+\lambda}{2(2\mu+\lambda)}\mathbf{x}\cdot\nabla G_{0}(\operatorname{div}(\rho\dot{\mathbf{u}}))+\frac{\mu+\lambda}{2(2\mu+\lambda)}\big[\operatorname{div}\mathbb{G}_{b}(\mathbf{x}F)-\mathbf{x}\cdot\nabla G_{b}(F)\big].$ (2.13)
We now analyze the last boundary integral terms in the equality above. In fact, $\mathbb{G}(\mathbf{x},\mathbf{y})=\widetilde{G}(\mathbf{x},\mathbf{y})\mathbb{I}$, where $\mathbb{I}=\{\delta_{ij}\}$ is the $2\times 2$ identity matrix and
$\widetilde{G}(\mathbf{x},\mathbf{y})=\frac{1}{2\pi}\big(\log|\mathbf{x}-\mathbf{y}|-\log|\mathbf{x}-\mathbf{y}^{*}|\big),$
with $\mathbf{y}^{*}=(y_{1},-y_{2})$. A straightforward computation yields the Poisson kernel
$\frac{\partial\widetilde{G}(\mathbf{x},\mathbf{y})}{\partial\mathbf{n}_{\mathbf{y}}}=\frac{x_{2}}{\pi|\mathbf{x}-\mathbf{y}|^{2}},\ \ \text{for}\ \mathbf{x}\in\mathbb{R}^{2}_{+},\ \mathbf{y}\in\partial\mathbb{R}^{2}_{+},$
and hence [^3] This Euler-type identity indicates that the boundary Poisson kernel is homogeneous of degree 1-1 in the variable $\mathbf{x}-\mathbf{y}$ (for $\mathbf{y}\in\partial\mathbb{R}^{2}_{+}$), reflecting the homogeneous scaling and the tangential translation invariance of the Green function for the half-plane.
$(\mathbf{y}-\mathbf{x})\cdot\frac{\partial\nabla_{\mathbf{x}}\widetilde{G}(\mathbf{x},\mathbf{y})}{\partial\mathbf{n}_{\mathbf{y}}}\Big|_{\mathbf{y}\in\partial\mathbb{R}^{2}_{+}}=\frac{\partial\widetilde{G}(\mathbf{x},\mathbf{y})}{\partial\mathbf{n}_{\mathbf{y}}}\Big|_{\mathbf{y}\in\partial\mathbb{R}^{2}_{+}}.$ (2.14)
This further implies that
$\operatorname{div}\mathbb{G}_{b}(\mathbf{x}F)-\mathbf{x}\cdot\nabla G_{b}(F)$ $=\operatorname{div}\int_{\partial\mathbb{R}^{2}_{+}}\frac{\partial\mathbb{G}(\mathbf{x},\mathbf{y})}{\partial\mathbf{n}_{\mathbf{y}}}\cdot\mathbf{y}F(\mathbf{y})\mathrm{d}s-\mathbf{x}\cdot\nabla\int_{\partial\mathbb{R}^{2}_{+}}\frac{\partial\widetilde{G}(\mathbf{x},\mathbf{y})}{\partial\mathbf{n}_{\mathbf{y}}}F(\mathbf{y})\mathrm{d}s$
$=\int_{\partial\mathbb{R}^{2}_{+}}\mathbf{y}\cdot\frac{\partial\nabla_{\mathbf{x}}\widetilde{G}(\mathbf{x},\mathbf{y})}{\partial\mathbf{n}_{\mathbf{y}}}F(\mathbf{y})\mathrm{d}s-\int_{\partial\mathbb{R}^{2}_{+}}\mathbf{x}\cdot\frac{\partial\nabla_{\mathbf{x}}\widetilde{G}(\mathbf{x},\mathbf{y})}{\partial\mathbf{n}_{\mathbf{y}}}F(\mathbf{y})\mathrm{d}s$
$=\int_{\partial\mathbb{R}^{2}_{+}}\frac{\partial\widetilde{G}(\mathbf{x},\mathbf{y})}{\partial\mathbf{n}_{\mathbf{y}}}F(\mathbf{y})\mathrm{d}s$
$=G_{b}(F)=F-G_{0}(\operatorname{div}(\rho\dot{\mathbf{u}})).$
Therefore, we obtain from (2.2) and (2.13) that
$\frac{3\mu+\lambda}{2(2\mu+\lambda)}F=$ $\frac{\mu+\lambda}{2(2\mu+\lambda)}\big[\operatorname{div}\mathbb{G}_{0}(\mathbf{x}\operatorname{div}(\rho\dot{\mathbf{u}}))-\mathbf{x}\cdot\nabla G_{0}(\operatorname{div}(\rho\dot{\mathbf{u}}))-G_{0}(\operatorname{div}(\rho\dot{\mathbf{u}}))\big]+\operatorname{div}\mathbb{G}_{0}(\rho\dot{\mathbf{u}})$
$+\frac{\mu}{2\mu+\lambda}[\operatorname{div}\mathbb{G}_{0}(\nabla P)-P].$
This boundary-term-free decomposition leads directly to the elliptic estimate below.
Lemma 2.7.
For any $1<q<\infty$, there exists a generic constant $\Lambda(q)>0$ depending only on $q$ such that
$\|\nabla\mathcal{Q}_{1}(\rho\dot{\mathbf{u}})\|_{L^{q}}+\|\nabla\mathcal{Q}_{2}(\rho\dot{\mathbf{u}})\|_{L^{q}}\leq\Lambda(q)\|\rho\dot{\mathbf{u}}\|_{L^{q}},$ (2.15)
$\|\mathcal{Q}_{3}(P)\|_{L^{q}}\leq\Lambda(q)\|P\|_{L^{q}}.$ (2.16)
Proof.
The proof follows from standard $L^{p}$ estimates, except for the following term in $\mathcal{Q}_{2}$:
$\operatorname{div}\mathbb{G}_{0}(\mathbf{x}\operatorname{div}(\rho\dot{\mathbf{u}}))-\mathbf{x}\cdot\nabla G_{0}(\operatorname{div}(\rho\dot{\mathbf{u}}))$ $=\operatorname{div}\int\mathbb{G}(\mathbf{x},\mathbf{y})\cdot\mathbf{y}\operatorname{div}_{\mathbf{y}}(\rho\dot{\mathbf{u}})\mathrm{d}\mathbf{y}-\mathbf{x}\cdot\nabla\int\widetilde{G}(\mathbf{x},\mathbf{y})\operatorname{div}_{\mathbf{y}}(\rho\dot{\mathbf{u}})\mathrm{d}\mathbf{y}$
$=\int(\mathbf{y}-\mathbf{x})\cdot\nabla_{\mathbf{x}}\widetilde{G}(\mathbf{x},\mathbf{y})\operatorname{div}_{\mathbf{y}}(\rho\dot{\mathbf{u}})\mathrm{d}\mathbf{y}.$ (2.17)
Note that the kernel
$\widehat{G}\triangleq(\mathbf{y}-\mathbf{x})\cdot\nabla_{\mathbf{x}}\widetilde{G}(\mathbf{x},\mathbf{y})=\frac{1}{2\pi}\frac{(\mathbf{x}-\mathbf{y})\cdot(\mathbf{x}-\mathbf{y}^{*})}{|\mathbf{x}-\mathbf{y}^{*}|^{2}}-\frac{1}{2\pi},\ \ \ \mathbf{y}^{*}=(y_{1},-y_{2})\neq\mathbf{y},$
is homogeneous of degree zero under the scaling vector $\mathbf{x}-\mathbf{y}$, and free of diagonal singularities as $\mathbf{x}\rightarrow\mathbf{y}$, since the reflected point $\mathbf{y}^{*}$ is separated from $\mathbf{x}$. This reflects a structural cancellation induced by the Dirichlet boundary condition, although the Green function itself possesses a logarithmic singularity. Finally, the boundedness of mixed derivatives $\nabla_{\mathbf{x}}\nabla_{\mathbf{y}}\widehat{G}$ combined with the standard $L^{p}$-framework yields (2.15). ∎
In addition, the bound of $\|\nabla\mathbf{u}\|_{L^{q}}$ is obtained via elliptic estimates for the constituent velocity fields, which arise from decomposing the source term of the linear Lamé system.
Lemma 2.8.
Let $(\rho,\mathbf{u})$ be a smooth solution to the problem (1.1)–(1.4). For any $1<q<\infty$, there exists a constant $C>0$ depending only on $\mu$, $\lambda$, and $q$ such that
$\|\nabla\mathbf{u}\|_{L^{q}}\leq C\|\sqrt{\rho}\|_{L^{q}}\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}+C\|P\|_{L^{q}}.$ (2.18)
Proof.
Let $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$ satisfy
$\begin{array}[]{ll}\begin{cases}\mu\Delta\mathbf{w}_{1}+(\mu+\lambda)\nabla\operatorname{div}\mathbf{w}_{1}=\rho\dot{\mathbf{u}},&\mathbf{x}\in\mathbb{R}^{2}_{+},\\ \mathbf{w}_{1}=0,&\mathbf{x}\in\partial\mathbb{R}^{2}_{+},\end{cases}\ \ \ \ \ \begin{cases}\mu\Delta\mathbf{w}_{2}+(\mu+\lambda)\nabla\operatorname{div}\mathbf{w}_{2}=\nabla P,&\mathbf{x}\in\mathbb{R}^{2}_{+},\\ \mathbf{w}_{2}=0,&\mathbf{x}\in\partial\mathbb{R}^{2}_{+}.\end{cases}\end{array}$ (2.20)
Accordingly, we set $\mathbf{u}=\mathbf{w}_{1}+\mathbf{w}_{2}$. It thus follows from the elliptic theory for the Lamé system that
$\|\nabla\mathbf{u}\|_{L^{q}}$ $\leq\|\nabla\mathbf{w}_{1}\|_{L^{q}}+\|\nabla\mathbf{w}_{2}\|_{L^{q}}$
$\leq C(q)\|\nabla^{2}\mathbf{w}_{1}\|_{L^{\frac{2q}{2+q}}}+\|\nabla\mathbf{w}_{2}\|_{L^{q}}$
$\leq C\|\rho\dot{\mathbf{u}}\|_{L^{\frac{2q}{2+q}}}+C\|P\|_{L^{q}}$ (2.21)
$\leq C\|\sqrt{\rho}\|_{L^{q}}\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}+C\|P\|_{L^{q}}.$
3 A priori estimates
In this section we derive necessary a priori bounds for the strong solution to (1.1)–(1.4), whose existence is guaranteed by Lemma 2.1. These bounds are independent of the lower bound of $\rho$, the initial regularity, and the time of existence. More precisely, fix $T>0$ and let $(\rho,\mathbf{u})$ be the corresponding strong solution on $\mathbb{R}^{2}_{+}\times(0,T]$.
First of all, if we multiply (1.1)1 by a cut-off function and use a standard limit procedure, one can obtain that
$\|\rho(t)\|_{L^{1}}=\|\rho_{0}\|_{L^{1}},\ \ 0\leq t\leq T.$ (3.1)
Next, for subsequent estimates, we define auxiliary functionals as
$A_{1}(T)\triangleq\sup_{t\in[0,T]}\big(\sigma\|\nabla\mathbf{u}\|_{L^{2}}^{2}\big)+\int_{0}^{T}\sigma\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}^{2}\mathrm{d}t,$ (3.2)
$A_{2}(T)\triangleq\sup_{t\in[0,T]}\big(\sigma^{3}\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}^{2}\big)+\int_{0}^{T}\sigma^{3}\|\nabla\dot{\mathbf{u}}\|_{L^{2}}^{2}\mathrm{d}t,$ (3.3)
$A_{3}(T)\triangleq\sup_{t\in[0,T]}\|\nabla\mathbf{u}\|_{L^{2}}^{2}+\int_{0}^{T}\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}^{2}\mathrm{d}t,$ (3.4)
where $\sigma=\sigma(t)\triangleq\min\{1,t\}$. Then we derive the following key a priori estimates on $(\rho,\mathbf{u})$.
Proposition 3.1.
Under the conditions of Theorem 1.1, there exist two positive constants $K\geq M^{2}$ and $\varepsilon$ depending only on $\alpha,\hat{\rho},M,a,\gamma,\mu,\lambda$, and $\eta_{0}$ such that if $(\rho,\mathbf{u})$ is a strong solution to the problem (1.1)–(1.4) satisfying
$\sup_{\mathbb{R}^{2}_{+}\times[0,T]}\|\rho\|_{L^{\theta}}\leq 2\hat{\rho},\ \ A_{1}(T)+A_{2}(T)\leq 2C_{0}^{\frac{5\alpha+4}{23\alpha+13}},\ \ A_{3}(\sigma(T))\leq 3K,$ (3.5)
then the following improved bounds hold
$\sup_{\mathbb{R}^{2}_{+}\times[0,T]}\|\rho\|_{L^{\theta}}\leq\frac{7}{4}\hat{\rho},\ \ A_{1}(T)+A_{2}(T)\leq C_{0}^{\frac{5\alpha+4}{23\alpha+13}},\ \ A_{3}(\sigma(T))\leq 2K,$ (3.6)
provided $C_{0}\leq\varepsilon$.
Before proving Proposition 3.1, we establish some necessary a priori estimates, see Lemmas 3.13.6 below. We begin with the basic energy estimate of $(\rho,\mathbf{u})$.
Lemma 3.1.
It holds that
$\sup_{0\leq t\leq T}\int\bigg(\frac{1}{2}\rho|\mathbf{u}|^{2}+\frac{P}{\gamma-1}\bigg)\mathrm{d}\mathbf{x}+\int_{0}^{T}\big[\mu\|\nabla\mathbf{u}\|_{L^{2}}^{2}+(\mu+\lambda)\|\operatorname{div}\mathbf{u}\|_{L^{2}}^{2}\big]\mathrm{d}t\leq C_{0}.$ (3.7)
Proof.
[…]
$\frac{P_{t}}{\gamma-1}+\frac{\operatorname{div}(P\mathbf{u})}{\gamma-1}+P\operatorname{div}\mathbf{u}=0.$ (3.8)
[…]
$\frac{\mathrm{d}}{\mathrm{d}t}\int\bigg(\frac{1}{2}\rho|\mathbf{u}|^{2}+\frac{P}{\gamma-1}\bigg)\mathrm{d}\mathbf{x}+\int\big[\mu|\nabla\mathbf{u}|^{2}+(\mu+\lambda)(\operatorname{div}\mathbf{u})^{2}\big]\mathrm{d}\mathbf{x}=0,$
which yields (3.7) after integrating the above equality with respect to $t$ over $(0,T)$. ∎
The next result concerns the preliminary bounds of $A_{1}(T)$ and $A_{2}(T)$.
Lemma 3.2.
Let $(\rho,\mathbf{u})$ be a smooth solution of (1.1)–(1.4) on $\mathbb{R}^{2}_{+}\times(0,T]$ satisfying $\|\rho\|_{L^{\theta}}\leq 2\hat{\rho}$. Then there exist positive constants $C$ and $\varepsilon_{1}$ depending only on $\alpha,\hat{\rho},a,\gamma,\mu$, and $\lambda$ such that
$A_{1}(T)\leq CC_{0}^{\frac{2(5\alpha+4)}{23\alpha+13}}+C\int_{0}^{T}\sigma\|\nabla\mathbf{u}\|_{L^{3}}^{3}\mathrm{d}t+C\int_{0}^{T}\sigma^{3}\|P\|_{L^{4}}^{4}\mathrm{d}t,$ (3.9)
$A_{2}(T)\leq CA_{1}(T)+C\Big(A_{1}(T)+A_{1}^{\frac{3}{2}}(T)\Big)A_{2}^{\frac{1}{2}}(T)+C\int_{0}^{T}\sigma^{3}\|P\|_{L^{4}}^{4}\mathrm{d}t$ (3.10)
provided $C_{0}\leq\varepsilon_{1}$.
Proof.
Using (3.7) and $P=a\rho^{\gamma}$, we obtain the following bound of $\|P\|_{L^{q}}$ with $q\in[1,60]$:
$\|P\|_{L^{q}}=a\|\rho\|_{L^{q\gamma}}^{\gamma}\leq a\|\rho\|_{L^{\gamma}}^{\frac{12\gamma(2\alpha+1)-q\gamma(\alpha-1)}{q(23\alpha+13)}}\|\rho\|_{L^{\theta}}^{\frac{12\gamma(q-1)(2\alpha+1)}{q(23\alpha+13)}}\leq C(a,\gamma,\hat{\rho})C_{0}^{\frac{12(2\alpha+1)-q(\alpha-1)}{q(23\alpha+13)}},\ \ 0\leq t\leq T.$ (3.11)
[…]
$\rho\dot{\mathbf{u}}+\nabla P=\mu\Delta\mathbf{u}+(\mu+\lambda)\nabla\operatorname{div}\mathbf{u}.$ (3.12)
Estimate for $A_{1}(T)$. Multiplying (3.12) by $\sigma\dot{\mathbf{u}}$ and integrating the resultant over $\mathbb{R}^{2}_{+}$, we get that
$\int\sigma\rho|\dot{\mathbf{u}}|^{2}\mathrm{d}\mathbf{x}=\int\sigma[-\dot{\mathbf{u}}\cdot\nabla P+\mu\Delta\mathbf{u}\cdot\dot{\mathbf{u}}+(\mu+\lambda)\dot{\mathbf{u}}\cdot\nabla\operatorname{div}\mathbf{u}]\mathrm{d}\mathbf{x}.$ (3.13)
[…]
$-\int\sigma\dot{\mathbf{u}}\cdot\nabla P\mathrm{d}\mathbf{x}$ $=\int\sigma P\operatorname{div}\mathbf{u}_{t}\mathrm{d}\mathbf{x}-\int\sigma\mathbf{u}\cdot\nabla\mathbf{u}\cdot\nabla P\mathrm{d}\mathbf{x}$
$=\frac{\mathrm{d}}{\mathrm{d}t}\int\sigma P\operatorname{div}\mathbf{u}\mathrm{d}\mathbf{x}-\int\sigma\operatorname{div}\mathbf{u}P^{\prime}(\rho)\rho_{t}\mathrm{d}\mathbf{x}-\int\sigma^{\prime}P\operatorname{div}\mathbf{u}\mathrm{d}\mathbf{x}-\int\sigma\mathbf{u}\cdot\nabla\mathbf{u}\cdot\nabla P\mathrm{d}\mathbf{x}$
$=\frac{\mathrm{d}}{\mathrm{d}t}\int\sigma P\operatorname{div}\mathbf{u}\mathrm{d}\mathbf{x}+\int\sigma(\operatorname{div}\mathbf{u})^{2}P^{\prime}(\rho)\rho\mathrm{d}\mathbf{x}+\int\sigma\mathbf{u}\cdot\nabla P\operatorname{div}\mathbf{u}\mathrm{d}\mathbf{x}-\int\sigma^{\prime}P\operatorname{div}\mathbf{u}\mathrm{d}\mathbf{x}$
$\quad-\int\sigma\mathbf{u}\cdot\nabla\mathbf{u}\cdot\nabla P\mathrm{d}\mathbf{x}$
$=\frac{\mathrm{d}}{\mathrm{d}t}\int\sigma P\operatorname{div}\mathbf{u}\mathrm{d}\mathbf{x}+\int\sigma(\operatorname{div}\mathbf{u})^{2}(P^{\prime}(\rho)\rho-P)\mathrm{d}\mathbf{x}+\int\sigma P\partial_{i}u^{j}\partial_{j}u^{i}\mathrm{d}\mathbf{x}-\int\sigma^{\prime}P\operatorname{div}\mathbf{u}\mathrm{d}\mathbf{x}$
$\leq\frac{\mathrm{d}}{\mathrm{d}t}\int\sigma P\operatorname{div}\mathbf{u}\mathrm{d}\mathbf{x}+C\int\sigma P|\nabla\mathbf{u}|^{2}\mathrm{d}\mathbf{x}+C\int\sigma^{\prime}P|\nabla\mathbf{u}|\mathrm{d}\mathbf{x}.$ (3.14)
Integrating the remaining terms in (3.13) by parts, we obtain that
$\mu\int\sigma\Delta\mathbf{u}\cdot\dot{\mathbf{u}}\mathrm{d}\mathbf{x}+(\mu+\lambda)\int\sigma\dot{\mathbf{u}}\cdot\nabla\operatorname{div}\mathbf{u}\mathrm{d}\mathbf{x}$
$=-\frac{\mu}{2}\frac{\mathrm{d}}{\mathrm{d}t}\big(\sigma\|\nabla\mathbf{u}\|^{2}_{L^{2}}\big)+\frac{\mu}{2}\sigma^{\prime}\|\nabla\mathbf{u}\|^{2}_{L^{2}}-\mu\int\sigma\partial_{i}u^{j}\partial_{i}(u^{k}\partial_{k}u^{j})\mathrm{d}\mathbf{x}-\frac{\mu+\lambda}{2}\frac{\mathrm{d}}{\mathrm{d}t}\big(\sigma\|\operatorname{div}\mathbf{u}\|^{2}_{L^{2}}\big)$
$\quad+\frac{\mu+\lambda}{2}\sigma^{\prime}\|\operatorname{div}\mathbf{u}\|^{2}_{L^{2}}-(\mu+\lambda)\int\sigma\operatorname{div}\mathbf{u}\operatorname{div}(\mathbf{u}\cdot\nabla\mathbf{u})\mathrm{d}\mathbf{x}$
$\leq-\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\big[\mu\sigma\|\nabla\mathbf{u}\|^{2}_{L^{2}}+(\mu+\lambda)\sigma\|\operatorname{div}\mathbf{u}\|^{2}_{L^{2}}\big]+C\|\nabla\mathbf{u}\|^{2}_{L^{2}}+C\int\sigma|\nabla\mathbf{u}|^{3}\mathrm{d}\mathbf{x},$ (3.15)
owing to $0\leq\sigma,\sigma^{\prime}\leq 1$. As a result, substituting (3) and (3) into (3.13) leads to
$\frac{\mathrm{d}}{\mathrm{d}t}\big(\sigma B(t)\big)+\int\sigma\rho|\dot{\mathbf{u}}|^{2}\mathrm{d}\mathbf{x}$ $\leq C\|\nabla\mathbf{u}\|^{2}_{L^{2}}+C\int\sigma|\nabla\mathbf{u}|^{3}\mathrm{d}\mathbf{x}+C\int\sigma^{\prime}P|\nabla\mathbf{u}|\mathrm{d}\mathbf{x}+C\int\sigma P|\nabla\mathbf{u}|^{2}\mathrm{d}\mathbf{x},$ (3.16)
where
$B(t)$ $\triangleq\frac{\mu}{2}\|\nabla\mathbf{u}\|^{2}_{L^{2}}+\frac{\mu+\lambda}{2}\|\operatorname{div}\mathbf{u}\|^{2}_{L^{2}}-\int P\operatorname{div}\mathbf{u}\mathrm{d}\mathbf{x}$
$\geq\frac{\mu}{2}\|\nabla\mathbf{u}\|^{2}_{L^{2}}+\frac{\mu+\lambda}{2}\|\operatorname{div}\mathbf{u}\|^{2}_{L^{2}}-\|P\|_{L^{2}}\|\operatorname{div}\mathbf{u}\|_{L^{2}}$
$\geq\frac{\mu}{2}\|\nabla\mathbf{u}\|^{2}_{L^{2}}+\frac{\mu+\lambda}{4}\|\operatorname{div}\mathbf{u}\|^{2}_{L^{2}}-C(\hat{\rho})C_{0}^{\frac{22\alpha+14}{23\alpha+13}},$ (3.17)
due to (3.11). Integrating (3.16) with respect to $t$ over $(0,T)$, one infers from (3.7) that
$\sup_{t\in[0,T]}\big(\sigma B(t)\big)+\int_{0}^{T}\int\sigma\rho|\dot{\mathbf{u}}|^{2}\mathrm{d}\mathbf{x}\mathrm{d}t$
$\leq CC_{0}+C\int_{0}^{T}\int\sigma|\nabla\mathbf{u}|^{3}\mathrm{d}\mathbf{x}\mathrm{d}t+C\int_{0}^{T}\int\sigma^{\prime}P|\nabla\mathbf{u}|\mathrm{d}\mathbf{x}\mathrm{d}t+C\int_{0}^{T}\int\sigma P|\nabla\mathbf{u}|^{2}\mathrm{d}\mathbf{x}\mathrm{d}t$
$\leq C(\hat{\rho})C_{0}^{\frac{11\alpha+7}{23\alpha+13}}+C\int_{0}^{T}\int\sigma|\nabla\mathbf{u}|^{3}\mathrm{d}\mathbf{x}\mathrm{d}t+C\int_{0}^{T}\int\sigma P|\nabla\mathbf{u}|^{2}\mathrm{d}\mathbf{x}\mathrm{d}t$ (3.18)
provided $C_{0}\leq 1$, where in the last inequality we have used the following estimate
$\int_{0}^{T}\int\sigma^{\prime}P|\nabla\mathbf{u}|\mathrm{d}\mathbf{x}\mathrm{d}t\leq\int_{0}^{\sigma(T)}\|P\|_{L^{2}}\|\nabla\mathbf{u}\|_{L^{2}}\mathrm{d}t\leq C(\hat{\rho})C_{0}^{\frac{11\alpha+7}{23\alpha+13}}\int_{0}^{\sigma(T)}\big(\|\nabla\mathbf{u}\|_{L^{2}}^{2}\big)^{\frac{1}{2}}\mathrm{d}t\leq C(\hat{\rho})C_{0}^{\frac{11\alpha+7}{23\alpha+13}}.$
To handle the last term in (3), we decompose the time axis into two parts: a short-time interval and a long one. If $C_{0}\leq 1$, it follows from (3.11), Lemma 2.8, and Hölder’s inequality that
$\int_{0}^{\sigma(T)}\int\sigma P|\nabla\mathbf{u}|^{2}\mathrm{d}\mathbf{x}\mathrm{d}t$ $\leq\int_{0}^{\sigma(T)}\sigma\|P\|_{L^{3}}\|\nabla\mathbf{u}\|_{L^{2}}\|\nabla\mathbf{u}\|_{L^{6}}\mathrm{d}t$
$\leq C(\hat{\rho})C_{0}^{\frac{7\alpha+5}{23\alpha+13}}\int_{0}^{\sigma(T)}\sigma\|\nabla\mathbf{u}\|_{L^{2}}\big(\|\sqrt{\rho}\|_{L^{6}}\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}+\|P\|_{L^{6}}\big)\mathrm{d}t$
$\leq C(\hat{\rho})C_{0}^{\frac{7\alpha+5}{23\alpha+13}}\int_{0}^{\sigma(T)}\sigma\big(\|\nabla\mathbf{u}\|_{L^{2}}^{2}+\|\rho\|_{L^{3}}\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}^{2}+\|P\|_{L^{6}}\|\nabla\mathbf{u}\|_{L^{2}}\big)\mathrm{d}t$
$\leq C(\hat{\rho})C_{0}^{\frac{2(5\alpha+4)}{23\alpha+13}}+C_{1}(\hat{\rho})C_{0}^{\frac{2(7\alpha+5)}{23\alpha+13}}A_{1}(T),$ (3.19)
$\int_{\sigma(T)}^{T}\int\sigma P|\nabla\mathbf{u}|^{2}\mathrm{d}\mathbf{x}\mathrm{d}t$ $\leq\int_{\sigma(T)}^{T}\|P\|_{L^{4}}\|\nabla\mathbf{u}\|_{L^{2}}\|\nabla\mathbf{u}\|_{L^{4}}\mathrm{d}t$
$\leq\int_{\sigma(T)}^{T}\|P\|_{L^{4}}\|\nabla\mathbf{u}\|_{L^{2}}\big(\|\sqrt{\rho}\|_{L^{4}}\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}+\|P\|_{L^{4}}\big)\mathrm{d}t$
$\leq C\int_{\sigma(T)}^{T}\big(\|\nabla\mathbf{u}\|_{L^{2}}^{2}+\|P\|_{L^{4}}^{2}\|\rho\|_{L^{2}}\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}^{2}+\|P\|_{L^{4}}^{4}\big)\mathrm{d}t$
$\leq CC_{0}+C_{2}(\hat{\rho})C_{0}^{\frac{3(7\alpha+5)}{23\alpha+13}}A_{1}(T)+\int_{\sigma(T)}^{T}\|P\|_{L^{4}}^{4}\mathrm{d}t.$ (3.20)
Putting (3) and (3.20) into (3) implies that
$\sup_{t\in[0,T]}\big(\sigma B(t)\big)+\int_{0}^{T}\int\sigma\rho|\dot{\mathbf{u}}|^{2}\mathrm{d}\mathbf{x}\mathrm{d}t$ $\leq C(\hat{\rho})C_{0}^{\frac{2(5\alpha+4)}{23\alpha+13}}+2C_{1}(\hat{\rho})C_{0}^{\frac{2(7\alpha+5)}{23\alpha+13}}A_{1}(T)$
$\quad+C\int_{0}^{T}\int\sigma|\nabla\mathbf{u}|^{3}\mathrm{d}\mathbf{x}\mathrm{d}t+\int_{\sigma(T)}^{T}\|P\|_{L^{4}}^{4}\mathrm{d}t,$
which along with (3.2) and (3) immediately yields (3.9) provided
$C_{0}\leq\varepsilon_{1}\triangleq\min\left\{1,\bigg(\frac{\mu}{8C_{1}(\hat{\rho})}\bigg)^{\frac{23\alpha+13}{2(7\alpha+5)}}\right\}.$
[…]
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\int\sigma^{3}\rho|\dot{\mathbf{u}}|^{2}\mathrm{d}\mathbf{x}-\frac{3}{2}\sigma^{2}\sigma^{\prime}\int\rho|\dot{\mathbf{u}}|^{2}\mathrm{d}\mathbf{x}$
$=-\int\sigma^{3}\dot{u}^{j}[\partial_{j}P_{t}+\operatorname{div}(\mathbf{u}\partial_{j}P)]\mathrm{d}\mathbf{x}+\mu\int\sigma^{3}\dot{u}^{j}\big[\Delta u_{t}^{j}+\operatorname{div}\big(\mathbf{u}\Delta u^{j}\big)\big]\mathrm{d}\mathbf{x}$
$\quad+(\mu+\lambda)\int\sigma^{3}\dot{u}^{j}[\partial_{j}\operatorname{div}\mathbf{u}_{t}+\operatorname{div}(\mathbf{u}\partial_{j}\operatorname{div}\mathbf{u})]\mathrm{d}\mathbf{x}\triangleq\sum_{i=1}^{3}\mathcal{I}_{i}.$ (3.21)
It follows from (3.8) and Cauchy–Schwarz inequality that
$\mathcal{I}_{1}$ $=\sigma^{3}\int P_{t}\operatorname{div}\dot{\mathbf{u}}\mathrm{d}\mathbf{x}-\sigma^{3}\int\dot{\mathbf{u}}\cdot\nabla\operatorname{div}(P\mathbf{u})\mathrm{d}\mathbf{x}+\sigma^{3}\int\dot{u}^{j}\operatorname{div}(P\partial_{j}\mathbf{u})\mathrm{d}\mathbf{x}$
$=\sigma^{3}\int\big(P_{t}+\operatorname{div}(P\mathbf{u})\big)\operatorname{div}\dot{\mathbf{u}}\mathrm{d}\mathbf{x}+\sigma^{3}\int\dot{\mathbf{u}}\cdot\nabla\mathbf{u}\cdot\nabla P\mathrm{d}\mathbf{x}+\sigma^{3}\int P\dot{\mathbf{u}}\cdot\nabla\operatorname{div}\mathbf{u}\mathrm{d}\mathbf{x}$
$=-\sigma^{3}\int(\gamma-1)P\operatorname{div}\mathbf{u}\operatorname{div}\dot{\mathbf{u}}\mathrm{d}\mathbf{x}-\sigma^{3}\int P\partial_{i}\dot{u}^{j}\partial_{j}u^{i}\mathrm{d}\mathbf{x}$
$\leq\frac{\mu}{8}\sigma^{3}\|\nabla\dot{\mathbf{u}}\|_{L^{2}}^{2}+C\sigma^{3}\|\nabla\mathbf{u}\|_{L^{4}}^{4}+C\sigma^{3}\|P\|_{L^{4}}^{4}.$ (3.22)
Using integration by parts, we have that
$\mathcal{I}_{2}$ $=\mu\sigma^{3}\int\dot{u}^{j}\big[\Delta\dot{u}^{j}-\Delta(\mathbf{u}\cdot\nabla u^{j})+\operatorname{div}\big(\mathbf{u}\Delta u^{j}\big)\big]\mathrm{d}\mathbf{x}$
$=\mu\sigma^{3}\int\big[-|\nabla\dot{\mathbf{u}}|^{2}+\dot{u}^{j}_{i}(u^{k}u_{k}^{j})_{i}-\dot{u}^{j}_{i}(u^{k}u^{j}_{i})_{k}-\dot{u}^{j}(u^{k}_{i}u^{j}_{i})_{k}\big]\mathrm{d}\mathbf{x}$
$=\mu\sigma^{3}\int\big[-|\nabla\dot{\mathbf{u}}|^{2}+\dot{u}^{j}_{i}(u^{k}u_{k}^{j})_{i}-\dot{u}^{j}_{i}(u^{k}u^{j}_{i})_{k}+\dot{u}_{k}^{j}(u^{k}_{i}u^{j}_{i})\big]\mathrm{d}\mathbf{x}$
$\leq-\frac{3\mu}{4}\sigma^{3}\|\nabla\dot{\mathbf{u}}\|_{L^{2}}^{2}+C\sigma^{3}\|\nabla\mathbf{u}\|_{L^{4}}^{4},$ (3.23)
$\mathcal{I}_{3}$ $=(\mu+\lambda)\sigma^{3}\int\dot{u}^{j}[\partial_{j}\operatorname{div}\mathbf{u}_{t}+\partial_{j}\operatorname{div}(\mathbf{u}\operatorname{div}\mathbf{u})-\operatorname{div}(\partial_{j}\mathbf{u}\operatorname{div}\mathbf{u})]\mathrm{d}\mathbf{x}$
$=-(\mu+\lambda)\sigma^{3}\int\operatorname{div}\dot{\mathbf{u}}[\operatorname{div}\mathbf{u}_{t}+\operatorname{div}(\mathbf{u}\operatorname{div}\mathbf{u})]\mathrm{d}\mathbf{x}-(\mu+\lambda)\sigma^{3}\int\dot{u}^{j}\operatorname{div}(\partial_{j}\mathbf{u}\operatorname{div}\mathbf{u})\mathrm{d}\mathbf{x}$
$=-(\mu+\lambda)\sigma^{3}\int\big[\operatorname{div}\dot{\mathbf{u}}\big(\operatorname{div}\dot{\mathbf{u}}-\partial_{i}u^{j}\partial_{j}u^{i}+(\operatorname{div}\mathbf{u})^{2}\big)-\partial_{j}\mathbf{u}\cdot\nabla\dot{u}^{j}\operatorname{div}\mathbf{u}\big]\mathrm{d}\mathbf{x}$
$\leq-\frac{\mu+\lambda}{2}\sigma^{3}\|\operatorname{div}\dot{\mathbf{u}}\|_{L^{2}}^{2}+\frac{\mu}{8}\sigma^{3}\|\nabla\dot{\mathbf{u}}\|_{L^{2}}^{2}+C\sigma^{3}\|\nabla\mathbf{u}\|_{L^{4}}^{4}.$ (3.24)
Thus, substituting (3)–(3.24) into (3), one finds that
$\frac{\mathrm{d}}{\mathrm{d}t}\int\sigma^{3}\rho|\dot{\mathbf{u}}|^{2}\mathrm{d}\mathbf{x}+\mu\sigma^{3}\|\nabla\dot{\mathbf{u}}\|_{L^{2}}^{2}+(\mu+\lambda)\sigma^{3}\|\operatorname{div}\dot{\mathbf{u}}\|_{L^{2}}^{2}\leq C\sigma^{2}\sigma^{\prime}\int\rho|\dot{\mathbf{u}}|^{2}\mathrm{d}\mathbf{x}+C\sigma^{3}\|P\|_{L^{4}}^{4}+C\sigma^{3}\|\nabla\mathbf{u}\|_{L^{4}}^{4}.$
Integrating the above inequality over $(0,T)$ implies that
$\sup_{t\in[0,T]}\int\sigma^{3}\rho|\dot{\mathbf{u}}|^{2}\mathrm{d}\mathbf{x}+\mu\int_{0}^{T}\sigma^{3}\|\nabla\dot{\mathbf{u}}\|_{L^{2}}^{2}\mathrm{d}t+(\mu+\lambda)\int_{0}^{T}\sigma^{3}\|\operatorname{div}\dot{\mathbf{u}}\|_{L^{2}}^{2}\mathrm{d}t$
$\leq CA_{1}(T)+C\int_{0}^{T}\sigma^{3}\|P\|_{L^{4}}^{4}\mathrm{d}t+C\int_{0}^{T}\sigma^{3}\|\nabla\mathbf{u}\|_{L^{4}}^{4}\mathrm{d}t.$ (3.25)
It remains to control the last term in (3). Following the decomposition of the velocity field as in Lemma 2.8, we have that
$\|\nabla\mathbf{u}\|_{L^{4}}^{4}$ $\leq C\|\nabla\mathbf{w}_{1}\|_{L^{4}}^{4}+C\|\nabla\mathbf{w}_{2}\|_{L^{4}}^{4}$
$\leq C\|\nabla\mathbf{w}_{1}\|_{L^{2}}\|\nabla\mathbf{w}_{1}\|_{L^{6}}^{3}+C\|P\|_{L^{4}}^{4}$
$\leq C(\|\nabla\mathbf{u}\|_{L^{2}}+\|\nabla\mathbf{w}_{2}\|_{L^{2}})\|\nabla^{2}\mathbf{w}_{1}\|_{L^{\frac{3}{2}}}^{3}+C\|P\|_{L^{4}}^{4}$
$\leq C(\|\nabla\mathbf{u}\|_{L^{2}}+\|P\|_{L^{2}})\|\rho\dot{\mathbf{u}}\|_{L^{\frac{3}{2}}}^{3}+C\|P\|_{L^{4}}^{4}$
$\leq C(\|\nabla\mathbf{u}\|_{L^{2}}+\|P\|_{L^{2}})\|\sqrt{\rho}\|_{L^{6}}^{3}\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}^{3}+C\|P\|_{L^{4}}^{4}$
$\leq C(\hat{\rho})(\|\nabla\mathbf{u}\|_{L^{2}}+1)\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}^{3}+C\|P\|_{L^{4}}^{4}$ (3.26)
provided $C_{0}\leq 1$. Hence, one deduces from (3.2) and (3.3) that
$\int_{0}^{T}\sigma^{3}\|\nabla\mathbf{u}\|_{L^{4}}^{4}\mathrm{d}t$ $\leq C(\hat{\rho})\sup_{t\in[0,T]}\big(\sigma^{\frac{1}{2}}\|\nabla\mathbf{u}\|_{L^{2}}\big)\sup_{t\in[0,T]}\big(\sigma^{\frac{3}{2}}\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}\big)\int_{0}^{T}\sigma\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}^{2}\mathrm{d}t$
$\quad+C(\hat{\rho})\sup_{t\in[0,T]}\big(\sigma^{\frac{3}{2}}\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}\big)\int_{0}^{T}\sigma^{\frac{3}{2}}\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}^{2}\mathrm{d}t+C\int_{0}^{T}\sigma^{3}\|P\|_{L^{4}}^{4}\mathrm{d}t$
$\leq C(\hat{\rho})A_{2}^{\frac{1}{2}}(T)\Big(A_{1}^{\frac{3}{2}}(T)+A_{1}(T)\Big)+C\int_{0}^{T}\sigma^{3}\|P\|_{L^{4}}^{4}\mathrm{d}t,$ (3.27)
which combined with (3) gives (3.10). ∎
On the basis of the preliminary bounds and the a priori hypothesis (3.5), we proceed to derive the key estimates for $A_{1}(T)$ and $A_{2}(T)$.
Lemma 3.3.
Under the assumption (3.5), there exists a positive constant $\varepsilon_{2}$ depending only on $\alpha,\hat{\rho},a,\gamma$, $\mu,\lambda$, and $K$ such that
$A_{1}(T)+A_{2}(T)\leq C_{0}^{\frac{5\alpha+4}{23\alpha+13}}$ (3.28)
provided $C_{0}\leq\varepsilon_{2}$ and $\frac{3\mu}{3\mu+\lambda}\leq\frac{1}{2\Lambda(4)}$.
Proof.
If $C_{0}\leq 1$, it follows from (3.5) and (3.7) that
$\int_{0}^{T}\|\nabla\mathbf{u}\|_{L^{2}}^{4}\mathrm{d}t$ $\leq\sup_{t\in[0,\sigma(T)]}\big(\|\nabla\mathbf{u}\|_{L^{2}}^{2}\big)\int_{0}^{\sigma(T)}\|\nabla\mathbf{u}\|_{L^{2}}^{2}\mathrm{d}t+\sup_{t\in[\sigma(T),T]}\big(\sigma\|\nabla\mathbf{u}\|_{L^{2}}^{2}\big)\int_{\sigma(T)}^{T}\|\nabla\mathbf{u}\|_{L^{2}}^{2}\mathrm{d}t$
$\leq CC_{0}A_{3}(\sigma(T))+CC_{0}A_{1}(T)$
$\leq C(K)C_{0}.$ (3.29)
From (3.5) and Lemma 3.2, we have that
$A_{1}(T)+A_{2}(T)\leq CC_{0}^{\frac{5\alpha+4}{23\alpha+13}}+C\int_{0}^{T}\sigma\|\nabla\mathbf{u}\|_{L^{3}}^{3}\mathrm{d}t+C\int_{0}^{T}\sigma^{3}\|P\|_{L^{4}}^{4}\mathrm{d}t$ (3.30)
provided $C_{0}\leq\varepsilon_{1}$. Using (3) and Lemma 2.8, we obtain that
$\int_{0}^{\sigma(T)}\sigma\|\nabla\mathbf{u}\|_{L^{3}}^{3}\mathrm{d}t$ $\leq\int_{0}^{\sigma(T)}\sigma\|\nabla\mathbf{u}\|_{L^{2}}^{\frac{3}{2}}\|\nabla\mathbf{u}\|_{L^{6}}^{\frac{3}{2}}\mathrm{d}t$
$\leq C(\hat{\rho})\int_{0}^{\sigma(T)}\sigma\|\nabla\mathbf{u}\|_{L^{2}}^{\frac{3}{2}}\big(\|\sqrt{\rho}\|_{L^{6}}\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}+\|P\|_{L^{6}}\big)^{\frac{3}{2}}\mathrm{d}t$
$\leq C(\hat{\rho})\sup_{t\in[0,\sigma(T)]}\big(\sigma^{\frac{1}{4}}\|\nabla\mathbf{u}\|_{L^{2}}^{\frac{1}{2}}\big)\int_{0}^{\sigma(T)}\big(\sigma\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}^{2}+\|\nabla\mathbf{u}\|_{L^{2}}^{4}\big)\mathrm{d}t$
$\quad+C(\hat{\rho})\sup_{t\in[0,\sigma(T)]}\|P\|_{L^{6}}^{\frac{3}{2}}\int_{0}^{\sigma(T)}(\|\nabla\mathbf{u}\|_{L^{2}}^{2})^{\frac{3}{4}}\mathrm{d}t$
$\leq C_{1}(\hat{\rho},K)A_{1}^{\frac{1}{4}}(T)(A_{1}(T)+C_{0})+C(\hat{\rho})C_{0}^{\frac{3}{4}}.$ (3.31)
Moreover, one gets from (3) that
$\int_{\sigma(T)}^{T}\sigma\|\nabla\mathbf{u}\|_{L^{3}}^{3}\mathrm{d}t$ $\leq\int_{\sigma(T)}^{T}\|\nabla\mathbf{u}\|_{L^{2}}\|\nabla\mathbf{u}\|_{L^{4}}^{2}\mathrm{d}t$
$\leq C\int_{\sigma(T)}^{T}\big(\|\nabla\mathbf{u}\|_{L^{2}}^{2}+\sigma^{3}\|\nabla\mathbf{u}\|_{L^{4}}^{4}\big)\mathrm{d}t$
$\leq CC_{0}+C(\hat{\rho})A_{2}^{\frac{1}{2}}(T)\Big(A_{1}^{\frac{3}{2}}(T)+A_{1}(T)\Big)+C\int_{0}^{T}\sigma^{3}\|P\|_{L^{4}}^{4}\mathrm{d}t,$
which combined with (3.5), (3.30), and (3) implies that
$A_{1}(T)+A_{2}(T)\leq 2C_{1}(\hat{\rho},K)C_{0}^{\frac{5(5\alpha+4)}{4(23\alpha+13)}}+C\int_{0}^{T}\sigma^{3}\|P\|_{L^{4}}^{4}\mathrm{d}t.$ (3.32)
We now turn to estimate the pressure term in (3.32). Recall the decomposition of $F$ in (2.9):
$(2\mu+\lambda)\operatorname{div}\mathbf{u}-P=F=\widetilde{F}+\frac{2\mu}{3\mu+\lambda}\mathcal{Q}_{3}=\Big[\mathcal{Q}_{1}(\rho\dot{\mathbf{u}})+\frac{\mu+\lambda}{3\mu+\lambda}\mathcal{Q}_{2}(\rho\dot{\mathbf{u}})\Big]+\frac{2\mu}{3\mu+\lambda}\mathcal{Q}_{3}(P).$ (3.33)
To proceed, we employ the equation (3.8):
$P_{t}+\operatorname{div}(P\mathbf{u})+(\gamma-1)P\operatorname{div}\mathbf{u}=0.$
After multiplying it by $3\sigma^{3}P^{2}$ and integration by parts, we obtain from Young’s inequality that
$\frac{3\gamma-1}{2\mu+\lambda}\sigma^{3}\|P\|_{L^{4}}^{4}$ $=-\frac{\mathrm{d}}{\mathrm{d}t}\int\sigma^{3}P^{3}\mathrm{d}\mathbf{x}+3\sigma^{2}\sigma^{\prime}\|P\|_{L^{3}}^{3}-\frac{3\gamma-1}{2\mu+\lambda}\int\sigma^{3}FP^{3}\mathrm{d}\mathbf{x}$
$\leq-\frac{\mathrm{d}}{\mathrm{d}t}\int\sigma^{3}P^{3}\mathrm{d}\mathbf{x}+3\sigma^{2}\sigma^{\prime}\|P\|_{L^{3}}^{3}+\frac{3(3\gamma-1)}{4(2\mu+\lambda)}\sigma^{3}\|P\|_{L^{4}}^{4}+\frac{3\gamma-1}{4(2\mu+\lambda)}\sigma^{3}\|F\|_{L^{4}}^{4}.$
Integrating the above inequality over $(0,T)$ gives that
$\int_{0}^{T}\sigma^{3}\|P\|_{L^{4}}^{4}\mathrm{d}t$ $\leq C\sup_{t\in[0,T]}\|P\|_{L^{3}}^{3}+C\int_{0}^{\sigma(T)}\|P\|_{L^{3}}^{3}\mathrm{d}t+\int_{0}^{T}\sigma^{3}\|F\|_{L^{4}}^{4}\mathrm{d}t$
$\leq CC_{0}^{\frac{3(7\alpha+5)}{23\alpha+13}}+8\int_{0}^{T}\sigma^{3}\|\widetilde{F}\|_{L^{4}}^{4}\mathrm{d}t+8\left(\frac{2\mu}{3\mu+\lambda}\right)^{4}\int_{0}^{T}\sigma^{3}\|\mathcal{Q}_{3}(P)\|_{L^{4}}^{4}\mathrm{d}t$
$\leq CC_{0}^{\frac{3(7\alpha+5)}{23\alpha+13}}+8\int_{0}^{T}\sigma^{3}\|\widetilde{F}\|_{L^{4}}^{4}\mathrm{d}t+8\left(\frac{2\mu\Lambda(4)}{3\mu+\lambda}\right)^{4}\int_{0}^{T}\sigma^{3}\|P\|_{L^{4}}^{4}\mathrm{d}t$
due to (3.33), (3.11), and (2.16). Then it follows from (2.15) that
$\int_{0}^{T}\sigma^{3}\|P\|_{L^{4}}^{4}\mathrm{d}t$ $\leq CC_{0}^{\frac{3(7\alpha+5)}{23\alpha+13}}+C\int_{0}^{T}\sigma^{3}\|\widetilde{F}\|_{L^{4}}^{4}\mathrm{d}t$
$\leq CC_{0}^{\frac{3(7\alpha+5)}{23\alpha+13}}+C\int_{0}^{T}\sigma^{3}\|\widetilde{F}\|_{L^{2}}\|\widetilde{F}\|_{L^{6}}^{3}\mathrm{d}t$
$\leq CC_{0}^{\frac{3(7\alpha+5)}{23\alpha+13}}+C\int_{0}^{T}\sigma^{3}\|\widetilde{F}\|_{L^{2}}\Big(\|\nabla\mathcal{Q}_{1}(\rho\dot{\mathbf{u}})\|_{L^{\frac{3}{2}}}^{3}+\|\nabla\mathcal{Q}_{2}(\rho\dot{\mathbf{u}})\|_{L^{\frac{3}{2}}}^{3}\Big)\mathrm{d}t$
$\leq CC_{0}^{\frac{3(7\alpha+5)}{23\alpha+13}}+C(\hat{\rho})\int_{0}^{T}\sigma^{3}\big(\|\nabla\mathbf{u}\|_{L^{2}}+\|P\|_{L^{2}}\big)\|\sqrt{\rho}\|_{L^{6}}^{3}\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}^{3}\mathrm{d}t$
$\leq CC_{0}^{\frac{3(7\alpha+5)}{23\alpha+13}}+C(\hat{\rho})\sup_{t\in[0,T]}\big(\sigma^{\frac{1}{2}}\|\nabla\mathbf{u}\|_{L^{2}}+\|P\|_{L^{2}}\big)\sup_{t\in[0,T]}\big(\sigma^{\frac{3}{2}}\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}\big)\int_{0}^{T}\sigma\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}^{2}\mathrm{d}t$
$\leq CC_{0}^{\frac{3(7\alpha+5)}{23\alpha+13}}+C_{3}(\hat{\rho})\big(A_{1}^{\frac{1}{2}}(T)+1\big)A_{1}(T)A_{2}^{\frac{1}{2}}(T)$ (3.34)
provided $\frac{3\mu}{3\mu+\lambda}\leq\frac{1}{2\Lambda(4)}$, where in the fourth inequality we have used
$\|\widetilde{F}\|_{L^{2}}=\|(2\mu+\lambda)\operatorname{div}\mathbf{u}-P-\frac{2\mu}{3\mu+\lambda}\mathcal{Q}_{3}\|_{L^{2}}\leq C\big(\|\nabla\mathbf{u}\|_{L^{2}}+\|P\|_{L^{2}}\big),$ (3.35)
$\|\nabla\widetilde{F}\|_{L^{\frac{3}{2}}}\leq C\|\nabla\mathcal{Q}_{1}(\rho\dot{\mathbf{u}})\|_{L^{\frac{3}{2}}}+C\|\nabla\mathcal{Q}_{2}(\rho\dot{\mathbf{u}})\|_{L^{\frac{3}{2}}}\leq C\|\rho\dot{\mathbf{u}}\|_{L^{\frac{3}{2}}}\leq C\|\sqrt{\rho}\|_{L^{6}}\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}.$ (3.36)
As a consequence, one deduces from (3.5), (3.32), and (3) that
$A_{1}(T)+A_{2}(T)$ $\leq 2C_{1}(\hat{\rho},K)C_{0}^{\frac{5(5\alpha+4)}{4(23\alpha+13)}}+C_{3}(\hat{\rho})\big(A_{1}^{\frac{1}{2}}(T)+1\big)A_{1}(T)A_{2}^{\frac{1}{2}}(T)$
$\leq 2C_{1}(\hat{\rho},K)C_{0}^{\frac{5\alpha+4}{4(23\alpha+13)}}C_{0}^{\frac{5\alpha+4}{23\alpha+13}}+2C_{3}(\hat{\rho})C_{0}^{\frac{5\alpha+4}{2(23\alpha+13)}}C_{0}^{\frac{5\alpha+4}{23\alpha+13}},$
which yields (3.28) as long as
$C_{0}\leq\varepsilon_{2}\triangleq\min\left\{\varepsilon_{1},\bigg(\frac{1}{4C_{1}(\hat{\rho},K)}\bigg)^{\frac{4(23\alpha+13)}{5\alpha+4}},\bigg(\frac{1}{4C_{3}(\hat{\rho})}\bigg)^{\frac{2(23\alpha+13)}{5\alpha+4}}\right\}.$
Next, we bound the auxiliary functional $A_{3}(T)$ and $\sigma\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}^{2}$ for small time.
Lemma 3.4.
Under the assumption (3.5), there exist positive constants $K\geq M^{2}$, $C(K)$, and $\varepsilon_{3}$ depending only on $\alpha,\hat{\rho},M,a,\gamma,\mu$, and $\lambda$ such that
$A_{3}(\sigma(T))=\sup_{t\in[0,\sigma(T)]}\|\nabla\mathbf{u}\|_{L^{2}}^{2}+\int_{0}^{\sigma(T)}\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}^{2}\mathrm{d}t\leq 2K,$ (3.37)
$\sup_{t\in[0,\sigma(T)]}\big(\sigma\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}^{2}\big)+\int_{0}^{\sigma(T)}\sigma\|\nabla\dot{\mathbf{u}}\|_{L^{2}}^{2}\mathrm{d}t\leq C(K)$ (3.38)
provided $C_{0}\leq\varepsilon_{3}$.
Proof.
Multiplying (3.12) by $\dot{\mathbf{u}}$ and integrating the resultant over $\mathbb{R}^{2}_{+}$, one has that
$\int\rho|\dot{\mathbf{u}}|^{2}\mathrm{d}\mathbf{x}=\int[-\dot{\mathbf{u}}\cdot\nabla P+\mu\Delta\mathbf{u}\cdot\dot{\mathbf{u}}+(\mu+\lambda)\dot{\mathbf{u}}\cdot\nabla\operatorname{div}\mathbf{u}]\mathrm{d}\mathbf{x}.$ (3.39)
Following the same argument as in (3.13)–(3), with $\sigma$ replaced by $1$, we obtain that
$\frac{\mathrm{d}}{\mathrm{d}t}B(t)+\int\rho|\dot{\mathbf{u}}|^{2}\mathrm{d}\mathbf{x}$ $\leq C\|\nabla\mathbf{u}\|^{2}_{L^{2}}+C\|\nabla\mathbf{u}\|^{3}_{L^{3}}+C\int P|\nabla\mathbf{u}|^{2}\mathrm{d}\mathbf{x},$ (3.40)
where $B(t)$ is given by (3). If $C_{0}\leq 1$, integrating (3.40) over $(0,\sigma(T))$ shows that
$\sup_{t\in[0,\sigma(T)]}B(t)+\int_{0}^{\sigma(T)}\int\rho|\dot{\mathbf{u}}|^{2}\mathrm{d}\mathbf{x}\mathrm{d}t\leq CM^{2}+C\int_{0}^{\sigma(T)}\|\nabla\mathbf{u}\|^{3}_{L^{3}}\mathrm{d}t+C\int_{0}^{\sigma(T)}\int P|\nabla\mathbf{u}|^{2}\mathrm{d}\mathbf{x}\mathrm{d}t$ (3.41)
due to (1.11) and (3.7).
It remains to estimate the last two terms in (3.41). From (3.4), (3.11), and Lemma 2.8, one sees that
$\int_{0}^{\sigma(T)}\|\nabla\mathbf{u}\|_{L^{3}}^{3}\mathrm{d}t$ $\leq\int_{0}^{\sigma(T)}\|\nabla\mathbf{u}\|_{L^{2}}^{\frac{3}{2}}\|\nabla\mathbf{u}\|_{L^{6}}^{\frac{3}{2}}\mathrm{d}t$
$\leq C\int_{0}^{\sigma(T)}\|\nabla\mathbf{u}\|_{L^{2}}^{\frac{3}{2}}\big(\|\sqrt{\rho}\|_{L^{6}}\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}+\|P\|_{L^{6}}\big)^{\frac{3}{2}}\mathrm{d}t$
$\leq C(\hat{\rho})\sup_{t\in[0,\sigma(T)]}\big(\|\nabla\mathbf{u}\|_{L^{2}}^{2}\big)^{\frac{3}{4}}\sup_{t\in[0,\sigma(T)]}\|\rho\|_{L^{3}}^{\frac{3}{4}}\int_{0}^{\sigma(T)}\big(\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}^{2}\big)^{\frac{3}{4}}\mathrm{d}t$
$\quad+C(\hat{\rho})\sup_{t\in[0,\sigma(T)]}\|P\|_{L^{6}}^{\frac{3}{2}}\int_{0}^{\sigma(T)}(\|\nabla\mathbf{u}\|_{L^{2}}^{2})^{\frac{3}{4}}\mathrm{d}t$
$\leq C_{4}(\hat{\rho})C_{0}^{\frac{3(7\alpha+5)}{4(23\alpha+13)}}A_{3}^{\frac{3}{2}}(\sigma(T))+C(\hat{\rho})C_{0}^{\frac{3}{4}},$ (3.42)
$\int_{0}^{\sigma(T)}\int P|\nabla\mathbf{u}|^{2}\mathrm{d}\mathbf{x}\mathrm{d}t$ $\leq\int_{0}^{\sigma(T)}\|P\|_{L^{3}}\|\nabla\mathbf{u}\|_{L^{2}}\|\nabla\mathbf{u}\|_{L^{6}}\mathrm{d}t$
$\leq C(\hat{\rho})C_{0}^{\frac{7\alpha+5}{23\alpha+13}}\int_{0}^{\sigma(T)}\|\nabla\mathbf{u}\|_{L^{2}}\big(\|\sqrt{\rho}\|_{L^{6}}\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}+\|P\|_{L^{6}}\big)\mathrm{d}t$
$\leq C(\hat{\rho})C_{0}^{\frac{3(7\alpha+5)}{2(23\alpha+13)}}A_{3}(\sigma(T))+C(\hat{\rho})C_{0}^{\frac{1}{2}}.$ (3.43)
Thus, it follows from (3.41)–(3.43) and (3.4) that
$A_{3}(\sigma(T))$ $\leq C(\hat{\rho},M)+C_{4}(\hat{\rho})C_{0}^{\frac{3(7\alpha+5)}{4(23\alpha+13)}}A_{3}^{\frac{3}{2}}(\sigma(T))+C(\hat{\rho})C_{0}^{\frac{3(7\alpha+5)}{2(23\alpha+13)}}A_{3}(\sigma(T))$
$\leq K+2C_{4}(\hat{\rho})C_{0}^{\frac{3(7\alpha+5)}{4(23\alpha+13)}}A_{3}^{\frac{3}{2}}(\sigma(T))$
for $K\triangleq C(\hat{\rho},M)$, which immediately yields (3.37) provided
$C_{0}\leq\varepsilon_{3}\triangleq\min\left\{\varepsilon_{2},\left(\frac{1}{18KC_{4}(\hat{\rho})}\right)^{\frac{4(23\alpha+13)}{3(7\alpha+5)}}\right\}.$
[…]
$\frac{\mathrm{d}}{\mathrm{d}t}\int\sigma\rho|\dot{\mathbf{u}}|^{2}\mathrm{d}\mathbf{x}+\mu\sigma\|\nabla\dot{\mathbf{u}}\|_{L^{2}}^{2}+(\mu+\lambda)\sigma\|\operatorname{div}\dot{\mathbf{u}}\|_{L^{2}}^{2}\leq C\sigma^{\prime}\int\rho|\dot{\mathbf{u}}|^{2}\mathrm{d}\mathbf{x}+C\sigma\|\nabla\mathbf{u}\|_{L^{4}}^{4}+C\sigma\|P\|_{L^{4}}^{4}.$
Integrating the above inequality over $(0,\sigma(T))$ implies that
$\sup_{t\in[0,\sigma(T)]}\int\sigma\rho|\dot{\mathbf{u}}|^{2}\mathrm{d}\mathbf{x}+\int_{0}^{\sigma(T)}\sigma\|\nabla\dot{\mathbf{u}}\|_{L^{2}}^{2}\mathrm{d}t$ $\leq CA_{3}(\sigma(T))+C\int_{0}^{\sigma(T)}\sigma\|\nabla\mathbf{u}\|_{L^{4}}^{4}\mathrm{d}t+C\int_{0}^{\sigma(T)}\sigma\|P\|_{L^{4}}^{4}\mathrm{d}t$
$\leq CK+C\int_{0}^{\sigma(T)}\sigma\|\nabla\mathbf{u}\|_{L^{4}}^{4}\mathrm{d}t+C\int_{0}^{\sigma(T)}\sigma\|P\|_{L^{4}}^{4}\mathrm{d}t.$ (3.44)
Arguing as in (3), (3), and (3), we obtain that
$\int_{0}^{\sigma(T)}\sigma\|\nabla\mathbf{u}\|_{L^{4}}^{4}\mathrm{d}t+\int_{0}^{\sigma(T)}\sigma\|P\|_{L^{4}}^{4}\mathrm{d}t$
$\leq CC_{0}^{\frac{3(7\alpha+5)}{23\alpha+13}}+C(\hat{\rho})\int_{0}^{\sigma(T)}\sigma(\|\nabla\mathbf{u}\|_{L^{2}}+1)\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}^{3}\mathrm{d}t+C(\hat{\rho})\int_{0}^{\sigma(T)}\sigma\big(\|\nabla\mathbf{u}\|_{L^{2}}+\|P\|_{L^{2}}\big)\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}^{3}\mathrm{d}t$
$\leq CC_{0}^{\frac{3(7\alpha+5)}{23\alpha+13}}+C(\hat{\rho})\sup_{t\in[0,\sigma(T)]}\big(\sigma^{\frac{1}{2}}\|\nabla\mathbf{u}\|_{L^{2}}+\|P\|_{L^{2}}+1\big)\sup_{t\in[0,\sigma(T)]}\big(\sigma^{\frac{1}{2}}\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}\big)\int_{0}^{\sigma(T)}\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}^{2}\mathrm{d}t$
$\leq CC_{0}^{\frac{3(7\alpha+5)}{23\alpha+13}}+C(\hat{\rho})A_{3}(\sigma(T))\sup_{t\in[0,\sigma(T)]}\big(\sigma\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}^{2}\big)^{\frac{1}{2}},$
which along with (3) and (3.37) yields (3.38). ∎
The following spatial weighted estimate plays an important role in bounding $\|\rho\|_{L^{\theta}}$.
Lemma 3.5.
Let $\bar{x}$ and $\alpha\in(1,2)$ be as in (1.8). There exists a positive constant $C$ depending only on $\alpha,\hat{\rho},M,a,\gamma,\mu,\lambda$, and $\eta_{0}$ such that, for any $0<t<T$,
$\sup_{s\in[0,t]}\|\bar{x}^{\alpha}\rho\|_{L^{1}}\leq C(1+t)^{4}.$ (3.45)
Proof.
[…]
$\frac{\mathrm{d}}{\mathrm{d}t}\int\rho(1+|\mathbf{x}|^{2})^{\frac{1}{2}}\mathrm{d}\mathbf{x}\leq C\int\rho|\mathbf{u}|\mathrm{d}\mathbf{x}\leq C\|\sqrt{\rho}\|_{L^{2}}\|\sqrt{\rho}\mathbf{u}\|_{L^{2}}.$
Integrating the above inequality with respect to time over $(0,t)$ together with (3.1) and (3.7) yields that
$\sup_{s\in[0,t]}\int\rho(1+|\mathbf{x}|^{2})^{\frac{1}{2}}\mathrm{d}\mathbf{x}\leq C(M)(1+t).$ (3.46)
For $N\geq 1$, let $\psi_{N}\in C^{\infty}_{0}(\widetilde{B}_{2N})$ satisfy
$\psi_{N}(\mathbf{x})=\begin{cases}1,\ \ |\mathbf{x}|\leq N,\\ 0,\ \ |\mathbf{x}|\geq 2N,\end{cases}\ \ \ 0\leq\psi_{N}\leq 1,\ \ \ |\nabla\psi_{N}|\leq CN^{-1}.$ (3.47)
Set
$\mathbf{y}(t)=\delta\mathbf{x}(1+t)^{-1}\log^{-\alpha}(e+t),$
[…]
$\frac{\mathrm{d}}{\mathrm{d}t}\int\rho\psi_{1}(\mathbf{y})\mathrm{d}\mathbf{x}$ $=\int\rho\mathbf{y}_{t}\cdot\nabla_{\mathbf{y}}\psi_{1}\mathrm{d}\mathbf{x}+\delta(1+t)^{-1}\log^{-\alpha}(e+t)\int\rho\mathbf{u}\cdot\nabla_{\mathbf{y}}\psi_{1}\mathrm{d}\mathbf{x}$
$\geq-\frac{C\delta}{(1+t)^{2}\log^{\alpha}(e+t)}\int\rho|\mathbf{x}|\mathrm{d}\mathbf{x}-\frac{C\delta}{(1+t)\log^{\alpha}(e+t)}$
$\geq-\frac{2C(M)\delta}{(1+t)\log^{\alpha}(e+t)}.$ (3.48)
Thus, one deduces from (3) and (1.10) that
$\int\rho\psi_{1}(\mathbf{y})\mathrm{d}\mathbf{x}\geq\int\rho_{0}\psi_{1}(\delta\mathbf{x})\mathrm{d}\mathbf{x}-C(\alpha,M)\delta\geq\int_{\widetilde{B}_{\eta_{0}}}\rho_{0}\mathrm{d}\mathbf{x}-C(\alpha,M)\delta\geq\frac{1}{4}$
by choosing $\delta\triangleq\big(\eta_{0}+4C(\alpha,M)\big)^{-1}$. Moreover, for $\eta_{1}\triangleq 2\delta^{-1}=2\eta_{0}+8C(\alpha,M)$, we derive that
$\inf_{0\leq t\leq T}\int_{\widetilde{B}_{\eta_{1}(1+t)\log^{\alpha}(e+t)}}\rho\mathrm{d}\mathbf{x}\geq\inf_{0\leq t\leq T}\int\rho\psi_{1}(\mathbf{y})\mathrm{d}\mathbf{x}\geq\frac{1}{4},$ (3.49)
as the desired (1.15).
Now we select the radius $\eta_{*}=\eta_{1}(1+t)\log^{\alpha}(e+t)$ in Lemma 2.4, which along with Lemma 2.3, (2.1), (3.1), (3.5), and (3.49) implies that
$\sup_{s\in[0,t]}\|\bar{x}^{-\varsigma}\mathbf{u}\|_{L^{r/\varsigma}}\leq C\sup_{s\in[0,t]}\eta_{*}^{2}(1+\|\rho\|_{L^{2}})\big(\|\sqrt{\rho}\mathbf{u}\|_{L^{2}}+\|\nabla\mathbf{u}\|_{L^{2}}\big)\leq C(\hat{\rho},M)(1+t)^{3-\frac{\alpha+1}{3}}$ (3.50)
[…]
$\frac{\mathrm{d}}{\mathrm{d}t}\int\bar{x}^{\alpha}\rho\mathrm{d}\mathbf{x}$ $\leq C\int\rho|\mathbf{u}|\bar{x}^{\alpha-1}\log^{2}(e+|\mathbf{x}|^{2})\mathrm{d}\mathbf{x}$
$\leq C\big\|\big(\rho(1+|\mathbf{x}|^{2})^{\frac{1}{2}}\big)^{\frac{\alpha+1}{3}}\big\|_{L^{\frac{3}{\alpha+1}}}\big\|\bar{x}^{-\frac{2-\alpha}{2}}\mathbf{u}\big\|_{L^{\frac{6}{2-\alpha}}}\big\|\rho^{\frac{2-\alpha}{3}}\big\|_{L^{\frac{6}{2-\alpha}}}$
$\quad\times\sup_{\mathbf{x}\in\mathbb{R}^{2}_{+}}\Big[(1+|\mathbf{x}|^{2})^{\frac{\alpha-2}{12}}\log^{\alpha+2}(e+|\mathbf{x}|^{2})\Big]$
$\leq C(1+t)^{3},$
which yields (3.45) after integrating the above inequality over $(0,t)$. ∎
Finally, by Zlotnik’s inequality, one can establish the desired uniform-in-time $L^{\theta}$ bound for the density.
Lemma 3.6.
Under the assumptions (1.12) and (3.5), there exists a positive constant $\varepsilon_{4}$ depending only on $\alpha,\hat{\rho},M,a,\gamma,\mu,\lambda$, and $\eta_{0}$ such that
$\sup_{t\in[0,T]}\|\rho\|_{L^{\theta}}\leq\frac{7}{4}\hat{\rho}$ (3.51)
provided $C_{0}\leq\varepsilon_{4}$.
Proof.
[…]
$\frac{\mathrm{d}}{\mathrm{d}t}\int\rho^{\theta}\mathrm{d}\mathbf{x}=-(\theta-1)\int\rho^{\theta}\operatorname{div}\mathbf{u}\mathrm{d}\mathbf{x}=-\frac{\theta-1}{2\mu+\lambda}\int\rho^{\theta}(P+F)\mathrm{d}\mathbf{x},$ (3.52)
which combined with (3.33) and Lemma 2.7 leads to
$\frac{\mathrm{d}}{\mathrm{d}t}\|\rho\|_{L^{\theta}}^{\theta}+\frac{\theta-1}{2\mu+\lambda}\|\rho\|_{L^{\theta+\gamma}}^{\theta+\gamma}$ $\leq\frac{\theta-1}{2\mu+\lambda}\int\rho^{\theta}|\widetilde{F}|\mathrm{d}\mathbf{x}+\frac{2\mu(\theta-1)}{(3\mu+\lambda)(2\mu+\lambda)}\int\rho^{\theta}\big|\mathcal{Q}_{3}(P)\big|\mathrm{d}\mathbf{x}$
$\leq\frac{\theta-1}{2\mu+\lambda}\|\rho\|_{L^{\theta}}^{\theta}\|\widetilde{F}\|_{L^{\infty}}+\frac{2\mu(\theta-1)}{(3\mu+\lambda)(2\mu+\lambda)}\|\rho^{\theta}\|_{L^{\frac{\theta+\gamma}{\theta}}}\|\mathcal{Q}_{3}(P)\|_{L^{\frac{\theta+\gamma}{\gamma}}}$
$\leq\frac{\theta-1}{2\mu+\lambda}\|\rho\|_{L^{\theta}}^{\theta}\|\widetilde{F}\|_{L^{\infty}}+\frac{2a\mu(\theta-1)}{(3\mu+\lambda)(2\mu+\lambda)}\Lambda\left(\frac{25\alpha+11}{\alpha-1}\right)\|\rho\|_{L^{\theta+\gamma}}^{\theta+\gamma}.$
Then it follows from (1.12) that
$\frac{\mathrm{d}}{\mathrm{d}t}\|\rho\|_{L^{\theta}}^{\theta}\leq-\frac{\theta-1}{3(2\mu+\lambda)}\|\rho\|_{L^{\theta+\gamma}}^{\theta+\gamma}+\frac{\theta-1}{2\mu+\lambda}\|\rho\|_{L^{\theta}}^{\theta}\|\widetilde{F}\|_{L^{\infty}}.$ (3.53)
Noting that
$\|\rho\|_{L^{\theta}}^{\theta}\leq\|\rho\|_{L^{\gamma}}^{\frac{\gamma^{2}}{\theta}}\|\rho\|_{L^{\theta+\gamma}}^{\frac{\theta^{2}-\gamma^{2}}{\theta}}\leq\big(a^{-1}(\gamma-1)C_{0}\big)^{\frac{\gamma}{\theta}}\|\rho\|_{L^{\theta+\gamma}}^{\frac{\theta^{2}-\gamma^{2}}{\theta}}\leq a^{-\frac{\gamma}{\theta}}(\gamma-1)^{\frac{\gamma}{\theta}}\|\rho\|_{L^{\theta+\gamma}}^{\frac{\theta^{2}-\gamma^{2}}{\theta}}$
provided $C_{0}\leq 1$, we infer from (3.53) that
$\frac{\mathrm{d}}{\mathrm{d}t}\|\rho\|_{L^{\theta}}^{\theta}\leq-\frac{\tilde{c}(\theta-1)}{3(2\mu+\lambda)}\big(\|\rho\|_{L^{\theta}}^{\theta}\big)^{\frac{\theta}{\theta-\gamma}}+\frac{\theta-1}{2\mu+\lambda}\|\rho\|_{L^{\theta}}^{\theta}\|\widetilde{F}\|_{L^{\infty}},$ (3.54)
where $\tilde{c}=\tilde{c}(\alpha,a,\gamma)=a^{-\frac{\theta^{2}}{\gamma(\theta-\gamma)}}(\gamma-1)^{\frac{\theta^{2}}{\gamma(\theta-\gamma)}}$.
Set
$y(t)=\|\rho\|_{L^{\theta}}^{\theta},\ \ f(y)=-\frac{\widetilde{c}(\theta-1)}{3(2\mu+\lambda)}y^{\frac{\theta}{\theta-\gamma}}(t),\ \ b(t)=\int_{0}^{t}\frac{\theta-1}{2\mu+\lambda}\|\widetilde{F}\|_{L^{\infty}}y(\tau)\mathrm{d}\tau,$
one deduces from (3.54) that
$y^{\prime}(t)\leq f(y)+b^{\prime}(t).$ (3.55)
Using Lemma 2.4 with $\eta_{*}=\eta_{1}(1+t)\log^{\alpha}(e+t)$, it follows from (2.1), (3.11), (3.45), and (3.49) that
$\|\rho\dot{\mathbf{u}}\|_{L^{4}(\mathbb{R}^{2}_{+})}\leq C(\hat{\rho},M)(1+t)^{4}\big(\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}(\mathbb{R}^{2}_{+})}+\|\nabla\dot{\mathbf{u}}\|_{L^{2}(\mathbb{R}^{2}_{+})}\big).$ (3.56)
This along with (3.35), (3.38), Lemma 2.2, and Lemma 2.7 yields that, for $0\leq t_{1}<t_{2}\leq\sigma(T)$,
$|b(t_{2})-b(t_{1})|$ $\leq C\int_{0}^{\sigma(T)}\|\widetilde{F}\|_{L^{\infty}}\mathrm{d}t$
$\leq C\int_{0}^{\sigma(T)}\|\widetilde{F}\|_{L^{2}}^{\frac{1}{3}}\|\nabla\widetilde{F}\|_{L^{4}}^{\frac{2}{3}}\mathrm{d}t$
$\leq C\int_{0}^{\sigma(T)}\big(\|\nabla\mathbf{u}\|_{L^{2}}+\|P\|_{L^{2}}\big)^{\frac{1}{3}}\|\rho\dot{\mathbf{u}}\|_{L^{4}}^{\frac{2}{3}}\mathrm{d}t$
$\leq C\int_{0}^{\sigma(T)}\big(\|\nabla\mathbf{u}\|_{L^{2}}+\|P\|_{L^{2}}\big)^{\frac{1}{3}}\Big(\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}^{\frac{2}{3}}+\|\nabla\dot{\mathbf{u}}\|_{L^{2}}^{\frac{2}{3}}\Big)\mathrm{d}t$
$\leq C\sup_{t\in[0,\sigma(T)]}\Big(\sigma^{\frac{1}{6}}\|\nabla\mathbf{u}\|_{L^{2}}^{\frac{1}{3}}\Big)\Bigg[\bigg(\int_{0}^{\sigma(T)}\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}^{2}\mathrm{d}t\bigg)^{\frac{1}{3}}\bigg(\int_{0}^{\sigma(T)}\sigma^{-\frac{1}{4}}\mathrm{d}t\bigg)^{\frac{2}{3}}$
$\quad+\bigg(\int_{0}^{\sigma(T)}\sigma\|\nabla\dot{\mathbf{u}}\|_{L^{2}}^{2}\mathrm{d}t\bigg)^{\frac{1}{3}}\bigg(\int_{0}^{\sigma(T)}\sigma^{-\frac{3}{4}}\mathrm{d}t\bigg)^{\frac{2}{3}}\Bigg]$
$\quad+C\sup_{t\in[0,\sigma(T)]}\|P\|_{L^{2}}^{\frac{1}{3}}\Bigg[\bigg(\int_{0}^{\sigma(T)}\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}^{2}\mathrm{d}t\bigg)^{\frac{1}{3}}\bigg(\int_{0}^{\sigma(T)}1\mathrm{d}t\bigg)^{\frac{2}{3}}$
$\quad+\bigg(\int_{0}^{\sigma(T)}\sigma\|\nabla\dot{\mathbf{u}}\|_{L^{2}}^{2}\mathrm{d}t\bigg)^{\frac{1}{3}}\bigg(\int_{0}^{\sigma(T)}\sigma^{-\frac{1}{2}}\mathrm{d}t\bigg)^{\frac{2}{3}}\Bigg]$
$\leq C_{1}(\hat{\rho},M,K)C_{0}^{\frac{5\alpha+4}{6(23\alpha+13)}}.$ (3.57)
Applying Lemma 2.5 with
$N_{1}=0,\ \ N_{0}=C_{1}(\hat{\rho},M,K)C_{0}^{\frac{5\alpha+4}{6(23\alpha+13)}},\ \ \xi^{*}=1,$
then it holds that
$f(\xi)=-\frac{\tilde{c}(\theta-1)}{3(2\mu+\lambda)}\xi^{\frac{\theta}{\theta-\gamma}}\leq-N_{1}=0,\ \ \ \text{for all}\ \xi\geq\xi^{*}=1.$
Moreover, one has that
$y(t)\leq\max\big\{\hat{\rho}^{\theta},1\big\}+N_{0}\leq\hat{\rho}^{\theta}+C_{1}(\hat{\rho},M,K)C_{0}^{\frac{5\alpha+4}{6(23\alpha+13)}}\leq\Big(\frac{3}{2}\hat{\rho}\Big)^{\theta}$ (3.58)
provided
$C_{0}\leq\varepsilon_{4,1}\triangleq\min\left\{\varepsilon_{3},\bigg(\frac{\hat{\rho}^{\theta}}{C_{1}(\hat{\rho},M,K)}\bigg)^{\frac{6(23\alpha+13)}{5\alpha+4}}\right\}.$
For $\sigma(T)\leq t_{1}<t_{2}\leq T$, we first derive large-time weighted estimates. Proceeding as in (3)–(3.24), but with $\sigma^{3}$ replaced by $t$, we obtain that
$\frac{\mathrm{d}}{\mathrm{d}t}\int t\rho|\dot{\mathbf{u}}|^{2}\mathrm{d}\mathbf{x}+\mu t\|\nabla\dot{\mathbf{u}}\|_{L^{2}}^{2}\leq C\int\rho|\dot{\mathbf{u}}|^{2}\mathrm{d}\mathbf{x}+Ct\|P\|_{L^{4}}^{4}+Ct\|\nabla\mathbf{u}\|_{L^{4}}^{4}.$
Integrating the above inequality over $(\sigma(T),T)$, one gets from (3) and (3) that
$\sup_{t\in[\sigma(T),T]}\big(t\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}^{2}\big)+\int_{\sigma(T)}^{T}t\|\nabla\dot{\mathbf{u}}\|_{L^{2}}^{2}\mathrm{d}t\leq C\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}^{2}+CT,$ (3.59)
which together with (3.56), Lemma 2.8, Young’s inequality, and (3.28) that
$|b(t_{2})-b(t_{1})|$ $\leq C\int_{t_{1}}^{t_{2}}\|\widetilde{F}\|_{L^{\infty}}\mathrm{d}t$
$\leq C\int_{t_{1}}^{t_{2}}\|\widetilde{F}\|_{L^{60}}^{\frac{15}{16}}\|\nabla\widetilde{F}\|_{L^{4}}^{\frac{1}{16}}\mathrm{d}t$
$\leq C\int_{t_{1}}^{t_{2}}\big(\|\sqrt{\rho}\|_{L^{60}}\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}+\|P\|_{L^{60}}\big)^{\frac{15}{16}}(1+t)^{\frac{1}{4}}\Big(\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}^{\frac{1}{16}}+\|\nabla\dot{\mathbf{u}}\|_{L^{2}}^{\frac{1}{16}}\Big)\mathrm{d}t$
$\leq C(\hat{\rho})\int_{\sigma(T)}^{T}\big(\sigma\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}^{2}+\sigma^{3}\|\nabla\dot{\mathbf{u}}\|_{L^{2}}^{2}\big)\mathrm{d}t+C\bigg(\int_{t_{1}}^{t_{2}}t\|\nabla\dot{\mathbf{u}}\|_{L^{2}}^{2}\mathrm{d}t\bigg)^{\frac{1}{4}}\bigg(\int_{t_{1}}^{t_{2}}1\mathrm{d}t\bigg)^{\frac{3}{4}}$
$\quad+C\sup_{t\in[t_{1},t_{2}]}\Big(t^{\frac{1}{4}}\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}^{\frac{1}{2}}\Big)\bigg(\int_{\sigma(T)}^{T}\|\sqrt{\rho}\dot{\mathbf{u}}\|_{L^{2}}^{2}\mathrm{d}t\bigg)^{\frac{1}{4}}\bigg(\int_{t_{1}}^{t_{2}}1\mathrm{d}t\bigg)^{\frac{3}{4}}+C(t_{2}-t_{1})$
$\leq\widetilde{C}(t_{2}-t_{1})+C_{5}(\hat{\rho})C_{0}^{\frac{5\alpha+4}{23\alpha+13}}.$ (3.60)
Selecting
$N_{0}=C_{5}(\hat{\rho})C_{0}^{\frac{5\alpha+4}{23\alpha+13}},\ \ N_{1}=\widetilde{C},\ \ \xi^{*}=\frac{3\widetilde{C}(2\mu+\lambda)}{\tilde{c}(\theta-1)}+1,$
it follows from Lemma 2.5 that
$f(\xi)=-\frac{\tilde{c}(\theta-1)}{3(2\mu+\lambda)}\xi^{\frac{\theta}{\theta-\gamma}}\leq-N_{1},\ \ \ \text{for all}\ \xi\geq\xi^{*}.$
Furthermore, it holds that
$y(t)\leq\max\big\{\hat{\rho}^{\theta},1\big\}+N_{0}\leq\hat{\rho}^{\theta}+C_{5}(\hat{\rho})C_{0}^{\frac{5\alpha+4}{23\alpha+13}}\leq\Big(\frac{3}{2}\hat{\rho}\Big)^{\theta}$ (3.61)
provided
$C_{0}\leq\varepsilon_{4,2}\triangleq\min\left\{\varepsilon_{3},\bigg(\frac{\hat{\rho}^{\theta}}{C_{5}(\hat{\rho})}\bigg)^{\frac{23\alpha+13}{5\alpha+4}}\right\}.$
Consequently, the desired (3.51) follows from (3.58) and (3.61) as long as $C_{0}\leq\varepsilon_{4}\triangleq\min\{\varepsilon_{4,1},\varepsilon_{4,2}\}$. ∎
Now we are ready to prove Proposition 3.1.
Proof of Proposition 3.1..
Proposition 3.1 follows from Lemmas 3.3, 3.4, and 3.6 if we select $\varepsilon=\varepsilon_{4}$. ∎
4 Proof of Theorem 1.1
With the a priori estimates established in Section 3, we are now in a position to prove Theorem 1.1.
Proof of Theorem 1.1..
Let $(\rho_{0},\mathbf{u}_{0})$ be initial data as described in the theorem. For $\epsilon>0$, let $j_{\epsilon}=j_{\epsilon}(\mathbf{x})$ be the standard mollifier, and define the approximate initial data $(\rho_{0}^{\epsilon},\mathbf{u}_{0}^{\epsilon})$:
$\rho_{0}^{\epsilon}$ $=[J_{\epsilon}\ast(\rho_{0}\mathbf{1}_{\mathbb{R}^{2}_{+}})]\mathbf{1}_{\mathbb{R}^{2}_{+}}+\phi\epsilon,\ \ \text{with}\ \phi=\phi(\epsilon,\mathbf{x})\triangleq\psi_{1/\epsilon}(\mathbf{x})+(1-\psi_{1/\epsilon}(\mathbf{x}))e^{-|\mathbf{x}|^{2}}\leq 1,$
where the cut-off function $\psi$ is given in (3.47), and $\mathbf{u}_{0}^{\epsilon}$ is the unique smooth solution to the elliptic equation
$\begin{cases}\Delta\mathbf{u}_{0}^{\epsilon}=\Delta(J_{\epsilon}\ast\mathbf{u}_{0}),&\mathbf{x}\in\mathbb{R}^{2}_{+},\\ \mathbf{u}_{0}^{\epsilon}=\mathbf{0},&x_{2}=0.\end{cases}$
Then we have
$\bar{x}^{\alpha}\rho_{0}^{\epsilon}\in L^{1},\ \ (\rho_{0}^{\epsilon}-\phi\epsilon)\in H^{2},\ \ \inf_{\mathbf{x}\in\widetilde{B}_{1/\epsilon}}\{\rho_{0}^{\epsilon}(\mathbf{x})\}\geq\epsilon,\ \ \mathbf{u}_{0}^{\epsilon}\in D^{2}\cap D_{0}^{1},\ \sqrt{\rho_{0}^{\epsilon}}\mathbf{u}_{0}^{\epsilon}\in L^{2},$
According to Lemma 2.1, there exists a time $T_{*}>0$ such that the problem (1.1)–(1.4) with initial data $(\rho_{0}^{\epsilon},\mathbf{u}_{0}^{\epsilon})$ admits a unique strong solution $(\rho^{\epsilon},{\bf u}^{\epsilon})$ on $\mathbb{R}^{2}_{+}\times(0,T_{*}]$ satisfying
$\bar{x}^{\alpha}\rho^{\epsilon}\in L^{\infty}(0,T;L^{1}),\ \ \rho^{\epsilon}-\phi\epsilon\in C([0,T_{*}];H^{2}),\ {\bf u}^{\epsilon}\in C([0,T_{*}];D^{2}\cap D^{1}),\ \inf\limits_{({\bf x},t)\in\widetilde{B}_{1/\epsilon}\times[0,T_{*}]}\rho^{\epsilon}({\bf x},t)>0.$
It follows from (1.7) and (3.2)–(3.4) that
$A_{1}(0)=A_{2}(0)=0,\ \ A_{3}(0)\leq M^{2}\leq K,$
which implies that there is $T_{1}\in(0,T_{*}]$ such that (3.5) holds for $T=T_{1}$. Set
[…]
Obviously $T^{*}\geq T_{1}>0$. We claim
$T^{*}=\infty.$ (4.2)
Otherwise, Lemma 3.5 gives that, for $0<T<T^{*}$,
$\sup_{t\in[0,T]}\|\bar{x}^{\alpha}\rho^{\epsilon}\|_{L^{1}}\leq C(T).$
Note that Lemmas 3.3, 3.4, and 3.6 hold independently of the lower bound of initial density, the time of existence, and the parameter $\epsilon$. From Proposition 3.1, one can deduce that the improved bounds in (3.6) are valid for all $0<T<T^{*}$ provided $C_{0}\leq\varepsilon$. These uniform estimates allow us to take the limit as $t\rightarrow T^{*}$. By the weak lower semicontinuity of the norms, we obtain a solution at time $T^{*}$ with regularity sufficient to serve as initial data. Applying Lemma 2.1 to this data yields an extension of solutions until some $T^{**}>T^{*}$ such that (3.5) holds for any $0<T<T^{**}$, which contradicts (4.1). Hence, (4.2) is true.
For any fixed $\tau$ and $T$ with $0<\tau<T<\infty$, it follows from Section 3 that the approximate solutions $(\rho^{\epsilon},\mathbf{u}^{\epsilon})$ satisfy the uniform bounds
$\begin{cases}\rho^{\epsilon}\in L^{\infty}(0,T;L^{1}\cap L^{\theta}),\ \ \sqrt{\rho^{\epsilon}}\mathbf{u}^{\epsilon}\in L^{\infty}(0,T;L^{2}),\\ \nabla\mathbf{u}^{\epsilon}\in L^{2}(0,T;L^{2}),\ \ \mathbf{u}^{\epsilon}\in L^{\infty}(\tau,T;H^{1}),\ \ \sqrt{\rho^{\epsilon}}\dot{\mathbf{u}}^{\epsilon}\in L^{\infty}(\tau,T;L^{2}),\end{cases}$
where $\theta>60\gamma$. Following an argument analogous to that in [22], we obtain that
$\lim_{\epsilon\rightarrow 0}\|\nabla\mathbf{u}_{0}^{\epsilon}-\nabla\mathbf{u}_{0}\|_{L^{2}}+\lim_{\epsilon\rightarrow 0}\|\sqrt{\rho_{0}^{\epsilon}}\mathbf{u}_{0}^{\epsilon}-\sqrt{\rho_{0}}\mathbf{u}_{0}\|_{L^{2}}=0.$ (4.3)
Moreover, the continuity equation implies that
$\partial_{t}\rho^{\epsilon}\ \ \text{is bounded in}\ L^{2}(\tau,T;H^{-1}).$
Hence, by the Aubin–Lions lemma, up to the extraction of a subsequence (still labeled by $\epsilon$), we have
$\rho^{\epsilon}-\phi\epsilon\rightarrow\rho\ \ \text{strongly in}\ C\big([0,T];L^{q}\big),\ \ \text{for any}\ q\in[1,\theta).$ (4.4)
Furthermore, the momentum equation yields that
$\partial_{t}(\rho^{\epsilon}\mathbf{u}^{\epsilon})\ \ \text{is bounded in}\ L^{2}\big(\tau,T;W^{-1,s}\big)\ \ \text{for some}\ s>1.$
This along with the uniform bound
$\rho^{\epsilon}\mathbf{u}^{\epsilon}\in L^{\infty}\big(0,T;L^{\frac{2\theta}{\theta+1}}\big)$
and the Aubin–Lions lemma implies that, up to a subsequence (still labeled by $\epsilon$),
$\rho^{\epsilon}\mathbf{u}^{\epsilon}\rightarrow\mathbf{m}\ \ \text{strongly in}\ C\big([\tau,T];L^{1+\zeta}_{\operatorname{loc}}\big)\ \ \text{for some small}\ \zeta.$ (4.5)
Define $\mathbf{u}=\mathbf{m}/\rho$ on $\{\rho>0\}$ and $\mathbf{u}=0$ on $\{\rho=0\}$. Finally, combining (4.4) and (4.5), by standard arguments (see [25, 30]) we infer that
$\nabla\mathbf{u}^{\epsilon}\rightarrow\nabla\mathbf{u}\ \ \text{strongly in}\ L^{2}(\tau,T;L^{2}(\{\rho>\delta\}))\ \ \text{for any}\ \delta>0.$
Consequently, passing to the limit $\epsilon\rightarrow 0$ shows that the limit $(\rho,\mathbf{u})$ is in fact a weak solution in the sense of Definition 1.1 satisfying (1.14) on $\mathbb{R}^{2}_{+}\times(0,T]$ for any $0<T<T^{*}=\infty$. ∎
Conflict of interest
The authors declare that they have no conflict of interest.
Data availability
No data was used for the research described in the article.